### v1

# A Self-Explainable Heterogeneous GNN for Relational Deep Learning

Abstract
Recently, significant attention has been given to the idea of viewing relational databases as heterogeneous graphs, enabling the application of graph neural network (GNN) technology for predictive tasks. However, existing GNN methods struggle with the complexity of the heterogeneous graphs induced by databases with numerous tables and relations. Traditional approaches either consider all possible relational meta-paths, thus failing to scale with the number of relations, or rely on domain experts to identify relevant meta-paths. A recent solution does manage to learn informative meta-paths without expert supervision, but assumes that a node’s class depends solely on the existence of a meta-path occurrence. In this work, we present a self-explainable heterogeneous GNN for relational data, that supports models in which class membership depends on aggregate information obtained from multiple occurrences of a meta-path. Experimental results show that in the context of relational databases, our approach effectively identifies informative meta-paths that faithfully capture the model’s reasoning mechanisms. It significantly outperforms existing methods in both synthetic and real-world scenarios.
1 Introduction
Graph Neural Networks (GNNs) have increasingly become the de-facto standard for many predictive tasks involving networked data, such as physical systems [1, 2], Knowledge Graphs [3] and social networks [4]. By learning effective node embeddings, GNNs offer a unified framework for addressing various graph-based tasks, including node classification, graph classification, and link prediction. However, similar to other representation learning paradigms, GNNs often exhibit a black-box nature in their predictions. Numerous solutions have been proposed for post-hoc explainability of GNN predictions, primarily at the instance-based level [5, 6, 7, 8]. Yet, as with other deep learning architectures, the ability of these approaches to genuinely reflect the underlying reasoning of the predictor has been called into question [9]. To tackle this challenge, self-explainable GNNs [10, 11, 12] have recently emerged, aiming to ensure that GNN predictions are grounded in interpretable elements, such as subgraphs [13, 14] or prototypes [15, 16].
Despite the widespread adoption of GNNs, most approaches are tailored for homogeneous graphs, where edge types are indistinguishable. While encoding edge types as features is a common workaround, the standard solution simply consists in concatenating the one-hot encoded edge type to node features, which eventually boils down to learning an edge-type specific bias. This limitation is particularly problematic in knowledge graphs, which typically feature numerous relations (corresponding to edge types), with only a few being pertinent to a specific predictive task—forming the so-called meta-paths. Existing approaches for heterogeneous GNNs either rely on domain experts to provide relevant meta-paths a priori [17, 18, 19, 20], or attempt to learn them from data by assigning different weights to various relations  [21, 22, 23, 24, 25, 26, 27, 28] , a solution that fails to scale with the number of candidate relations.
Recently, MP-GNN [29] has been proposed as a solution for learning meta-paths without requiring user supervision. This approach employs a scoring function to predict the potential informativeness of partial meta-paths, enabling efficient exploration of the combinatorial space of candidate meta-paths. However, a key limitation of MP-GNN is its assumption that a node’s class is primarily dependent on the existence of a meta-path instance. While this assumption may be reasonable for knowledge graphs, it is unrealistic when dealing with relational databases, where entities are characterized by numerous categorical and numerical attributes. On the other hand, this same complexity makes relational databases particularly well-suited for GNN technology, as evidenced by the growing interest in what is now being termed relational deep learning [30].
Figure 1: Left: Relational database schema for a medical domain. Right: Heterogeneous graph representation of (part of) the database. The highlighted subgraph shows a prototypical counts-of-counts pattern characterising positive patients, namely having at least two exempt prescriptions (indicated as true), each containing at least two medications. Existing heterogeneous GNNs struggle with these patterns as they need to learn a separate weight matrix for each edge type in the graph, while MPS-GNN is capable of learning them without any direct user supervision.
In this paper, we extend the concept behind MP-GNN beyond the simple existential quantification of meta-paths. We introduce Meta-Path Statistics GNN (MPS-GNN), an approach that automatically identifies relevant meta-paths, where the informative content is determined by learnable statistics computed on their realizations. These include counts-of-counts statistics such as having at least two exempt prescriptions with at least two medications each to characterize patients with severe illness, as shown in Figure 1. An experimental evaluation on both synthetic and real-world relational database tasks demonstrate the significant advantages of the proposed solution over existing alternatives.
Additionally, results show how the meta-path learning strategy behind MPS-GNN renders it inherently and genuinely self-explainable, in contrast to many existing self-explainable GNN architectures whose explanations often lack fidelity [11].
2 Related Work
Relational deep learning [30] has recently emerged as a paradigm advocating the application of deep learning technology, and GNNs in particular, to relational databases. The rationale behind this research direction is the popularity of relational databases as a mean to store relational information in a variety of application domains, combined with the fact that relational databases can be seen as heterogeneous graphs, with tables converted into sets of nodes and relations into (typed) edges between table entries [31]. This transformation allows the application of heterogeneous Graph Neural Networks (GNNs) to this kind of data.
A common characteristics of most heterogeneous graphs, including those deriving from knowledge graphs and relational databases, is that only few relations convey relevant information when targeting a specific predictive task. For this reason, plain GNNs, that do not distinguish between edge types, struggle with these type of graphs. The most popular line of research for heterogeneous GNNs identifies meta-paths, i.e., sequences of relations, as primary sources of information. Existing approaches to incorporate meta-paths follow in two main categories. The former requires domain experts to identify relevant meta-paths for the task at hand [17, 18, 19, 20]. Clearly, this approach is suboptimal as it requires this domain information to be readily available. The alternative solutions, either utilize relation-specific graph convolutions, capturing relational patterns with distinct parameters or dedicated components for each type of relation  [21, 22, 24, 25] or focuses on graph transformation and multi-view learning to enhance relational representations [23, 26, 27, 28]. Although these methods are effective with a limited number of relations, their performance quickly deteriorates as the number of candidate relations grows.
Recently, a novel approach named MP-GNN [29] has been introduced to tackle the aforementioned challenges and automatically learn relevant meta-paths from data. The approach leverages a scoring function predicting the potential informativeness of partial meta-paths to guide the search in the combinatorial space of candidate meta-paths. A major limitation of this approach is the fact that it assumes that the existential quantification of the meta-path is informative for the class label. This assumption makes the approach unsuitable for relational deep learning tasks, in which statistics extracted from table attributes are arguably crucial to characterize predictive targets. Our approach substantially generalizes the MP-GNN method, by designing a scoring function that can predict the informativeness of partial meta-paths in terms of the statistics that could be constructed on top of their realizations. This extension is crucial in allowing MPS-GNN to be effectively applied to relational deep learning settings, as shown by our experimental evaluation. More details about the comparison between MP-GNN and MPS-GNN in A.1.
Explainability in GNN is a major research trend, with plenty of approaches for post-hoc explanation of GNN predictions (instance-based explainability [5, 6, 7, 8]) and others aiming to explain the GNN model as a whole (model-based explainability [32, 33, 34, 35]). Specialized metrics have been developed to estimate the faithfulness of an explanation in relation to the method’s input processing behavior [36, 37, 38, 39, 40]. Following a similar evolution in the XAI literature of convolutional neural networks, self-explainable GNNs [13, 14, 15, 16] have recently been proposed to encourage GNN models to adhere to their explanations by design. However, experimental studies have questioned the faithfulness of the explanations provided by these approaches [11, 41], highlighting the difficulty of achieving genuine explainability with GNNs. A key advantage of our proposed method is that meta-paths can naturally serve as model-level explanations, making MPS-GNN the first truly self-explainable GNN designed for relational deep learning applications. Our experimental evaluation confirms the faithfulness of the explanations to the model’s behaviour.
3 Preliminaries
This section presents the core concepts that will be utilized in the rest of the paper.
Definition 3.1 (Relational Database).
A relational database $(\mathcal{T},\mathcal{L})$ consists of a collection of tables $\mathcal{T}=T_{1},...T_{n}$ and links between these tables $\mathcal{L}\subseteq T\times T$ Each table is a set $T=\{u_{1},...,u_{n}\}$ where the elements $u\in T$ are referred to as rows or entities. Each entity is a tuple $u=(\mathcal{P}_{u},\mathcal{K}_{u},a_{u})$ where $\mathcal{P}_{u}$ is the Primary Key that uniquely identifies the entity $u$; $\mathcal{K}_{u}$ is the set of Foreign Keys corresponding to a primary key in other tables, thus connecting the tables; $a_{u}$ corresponds to the Attributes of the entity $u$.
Definition 3.2 (Heterogeneous graph).
[…]
For a node $v$ and a relation $r$ we denote with $\mathcal{N}^{r}_{v}$ the set of nodes that can be reached from $v$ by following relation $r$. We refer to this set as $r$-neighbors.
Definition 3.3 (Meta-path).
A meta-path $mp$ is a sequence of relations defined on a heterogeneous graph $\cal G$, represented as $\xrightarrow{r_{1}}\xrightarrow{r_{2}}\ldots\xrightarrow{r_{L}}$, where $r_{1},\ldots,r_{L}$ are relations in $\mathcal{R}$. knowledge graphs.
From relational database to heterogeneous graph A relational database can be interpreted as an heterogeneous graph where row $u$ becomes node $v$; attributes $a_{u}$ become node features $x_{v}$; links $\mathcal{L}$ between entries of two tables are identified by the pair of primary $\mathcal{P}$ and foreign $\mathcal{K}$ keys in the two tables. Each pair of connected tables, originates a relation in the graph, specified by $r$.
4 Methodology
For now we restrict attention to binary node classification problems. Our main problem is to construct meta-paths $\xrightarrow{r_{1}}\xrightarrow{r_{2}}\ldots\xrightarrow{r_{L}}$ that are predictive features for the class label. When considering a meta-path as a feature, we are thinking of possible numerical features that can be defined by collecting and aggregating information that is found along all occurrences of the meta-path in a concrete data graph, such as the count-of-count feature illustrated in Figure 1. We construct meta-paths following a strategy that is
  • Greedy: a partially constructed meta-path $\xrightarrow{r_{1}}\xrightarrow{r_{2}}\ldots\xrightarrow{r_{i}}$ is extended by a next relation $\xrightarrow{r_{i+1}}$ without lookahead for possible completions $\xrightarrow{r_{i+1}}\ldots\xrightarrow{r_{L}}$. Similar as in [29], we try to estimate the potential informativeness of nodes reached by $\xrightarrow{r_{i+1}}$ by learned weights associated with the nodes. These weights represent putative features that can either be directly materialized as functions of the nodes’ attributes, or that can be constructed as features of meta-path extensions starting at the node.
  • Local: the meta-path construction step $\xrightarrow{r_{i}}\xrightarrow{r_{i+1}}$ is performed based only on local consideration of the nodes reached by $r_{i}$, and their $r_{i+1}$ successors. The already constructed meta-path prefix $\xrightarrow{r_{1}}\xrightarrow{r_{2}}\ldots\xrightarrow{r_{i}}$ plays no explicit role in this step. We realize this locality by defining for each step a surrogate classification task for the nodes reached by $r_{i}$. The problem of extending the constructed meta-path prefix then translates into the problem of finding the first relation for a relevant meta-path solving the surrogate problem. The surrogate problems take the form of weighted multi-instance classification tasks.
Note that both the greedy and local properties mirror core principles of growing decision trees, which are built incrementally, adding one relation at a time based on solving local classification sub-tasks.
4.1 Weighted multi-instance classification
[…]
Our goal is to predict the bag label via discriminant functions of the form
$F({\mathcal{B}})=\sum_{v\in{\mathcal{B}}}\alpha(v,{\mathcal{B}})\cdot F(v),$ (1)
where $F(v)$ is a learnable node feature function. Specifically, we consider functions that are parameterized by a relation $r$, and are of the form
$F(v,r,\Theta,{\boldsymbol{w}})=\Theta\cdot x_{v}\cdot\sum_{u\in\mathcal{N}^{r}_{v}}w_{u},$ (2)
where $x_{v}$ denotes the attribute vector of $v$, $\Theta$ is a trainable parameter vector, and ${\boldsymbol{w}}$ is a vector of trainable weights $w_{u}\in[0,1]$ assigned to $v$’s $r$-neighbors. The node feature function value for $v$ can, thus, be determined by $v$’s own attributes, or by aggregating the putative feature $w_{u}$ of its successors – or a combination of both elements.
Ideally, the discriminant function separates the classes in the sense that for any pair ${\mathcal{B}}^{+}\in{\mathcal{S}}^{+},{\mathcal{B}}^{-}\in{\mathcal{S}}^{-}$ we have $F({\mathcal{B}}^{+})>F({\mathcal{B}}^{-})$. We note that this problem would be trivially solvable e.g. in the case where every positive bag contains a node $v$ that has an $r$-successor, which is not also an $r$-successor of some node in a negative bag. Then assigning a weight of 1 to all such $r$-successor nodes, and a weight of 0 to all other nodes, would separate the classes. In reality, the complex connectivity of relations will preclude such simple solutions, and perfect separation in general. We therefore use as our learning objective the relaxed loss function
$L(r,\Theta,{\boldsymbol{w}})=\sum_{{\mathcal{B}}^{+}\in\mathcal{S}^{+},{\mathcal{B}}^{-}\in\mathcal{S}^{-}}\sigma\Big{(}F({\mathcal{B}}^{-},r,\Theta,{\boldsymbol{w}})-F({\mathcal{B}}^{+},r,\Theta,{\boldsymbol{w}})\Big{)},$ (3)
where $\sigma$ is the sigmoid function. In practice, we approximate (3) by a random sample of positive and negative bags.
4.2 Relation Scoring
The initial weighted multi-instance classification problem for our meta-path construction process is defined by letting each positive (negative) target node $v$ form a one-element bag $\{v\}$ with the corresponding label, and weight $\alpha(v,\{v\})=1$. Denote with ${\mathcal{S}}_{1}^{+},{\mathcal{S}}_{1}^{-}$ the sets of all initial positive and negative bags, respectively. At all iterations we select the relation that minimizes the loss
$L(r)=\min_{\Theta,{\boldsymbol{w}}}L(r,\Theta,{\boldsymbol{w}}).$ (4)
If for some relation $r$ this loss is effectively minimized by suitable parameters $\Theta$, such that the $\sum w_{u}$ term in (2) plays no significant role, then the meta-path construction can terminate at this point (without adding $r$ to the meta-path). If, on the other hand, the $r$-successors and their weights ${\boldsymbol{w}}$ are needed to solve the current multi-instance classification task, then we need to extend the meta-path by $r$, and we need to capture the putative node features ${\boldsymbol{w}}$ by an actual node feature. A possible approach would be to set this up as a node regression task with target values $w_{u}$. However, due to the often very large set of alternative optimal solutions ${\boldsymbol{w}}$ in the minimization (4), this would lead to a too restrictive task. Our goal is to approximate the whole space of possible regression tasks defined by alternative ${\boldsymbol{w}}$ as a single weighted multi-instance classification task. For this, let $r_{i}$ denote the relation found to minimize (4) in iteration $i$ with optimal parameters $\Theta_{i}$. For each positive bag ${\mathcal{B}}^{+}_{i}\in\mathcal{S}_{i}^{+}$ define the new bag
${\mathcal{B}}^{+}_{i+1}=\cup_{v\in{\mathcal{B}}^{+}_{i}}\mathcal{N}^{r_{i}}_{v}$ (5)
containing the $r_{i}$-children of the nodes in ${\mathcal{B}}^{+}_{i}$. Similarly for negative bags. For $u\in{\mathcal{B}}^{+}_{i+1}$ define the node weight
$\alpha(u,{\mathcal{B}}^{+}_{i+1})=\sum_{v\,:\,u\in\mathcal{N}^{r_{i}}_{v}}\Theta_{i}\cdot x_{v}\cdot\alpha(v,{\mathcal{B}}^{+}_{i}).$ (6)
Thus, nodes have a bigger weight in the subsequent multi-instance classification task, if they are $r_{i}$- successors of many nodes from the previous task, that make a significant contribution to the discriminant function (1) learned for the previous task. The sets of all ${\mathcal{B}}^{+}_{i+1}({\mathcal{B}}^{-}_{i+1})$ form the new training sets ${\mathcal{S}}^{+}_{i+1}({\mathcal{S}}^{-}_{i+1})$.
[…]
Figure 2: Top: First iteration scoring function. For simplicity, we assume the features of the patient nodes to be uninformative, represented by a constant ’$T$’(rue) vector. Top Left: By simultaneous optimization of $w_{2},\ldots,w_{6}$ and $\Theta$ parameters, the loss for relation ’a’ can be brought arbitrarily close to 0 by choosing $\Theta$ so that $Z:=\Theta\cdot T\gg 0$. Top Right:The loss for relation "b" is constant $1/2$ for any choice of the parameters $w_{7},\Theta$. Bottom: Second iteration of the scoring function. With $Z$ as in first iteration we have $\alpha(2,\mathcal{B}_{2})=\alpha(3,\mathcal{B}_{2})=\alpha(5,\mathcal{B}_{3})=\alpha(6,\mathcal{B}_{3})=Z$ and $\alpha(4,\mathcal{B}_{2})=\alpha(4,\mathcal{B}_{3})=2Z$ (details in App. A.2). Bottom Left: With relation "c" optimal parameters give 0 loss. Bottom Right: Relation "d" obtains non-optimal loss (1/2).
4.3 MPS-GNN
Similarly to MP-GNN, in the MPS-GNN framework a meta-path $r_{1},...,r_{L}$ defines a multi-relational GNN with $L$ layers. In this setup, each layer of the network corresponds to a specific relation in the meta-path: the initial layer is linked to the final relation $r_{L}$, progressing sequentially until the last layer, which corresponds to $r_{1}$. Our forward model then takes the form:
$h^{(l+1)}_{v}=\sigma\Big{(}W_{0}^{(l)}h^{(l)}_{v}+W_{neigh}^{(l)}\sum_{u\in\mathcal{N}_{v}^{r_{L-l}}}h^{(l)}_{u}+W_{1}^{(l)}h^{(0)}_{v}\Big{)}$ (7)
where $\mathcal{N}_{v}^{r_{L-l}}$ are neighbours of node $v$ under relation $r_{L-l}$, $h^{(l)}_{v}$ is the embedding of node $v$ in layer $l$, $h^{(0)}_{v}=x_{v}$ is the feature vector of node $v$, while $W_{0}^{(l)}$, $W_{neigh}^{(l)}$ and $W_{1}^{(l)}$ are learnable parameter vectors. Note that the latter term $W_{1}^{(l)}h^{(0)}_{v}$, which is missing the original MP-GNN [29], represents a skip connection between the input and the $l+1$ layer. This allows the network to access the node attributes at each layer, which is essential for enabling the MPS-GNN to capture node features corresponding to the $\Theta\cdot x_{v}$ terms in the meta-path construction as shown in the ablation study in A.3). As in [29], the definition can be generalized to multiple meta-paths by concatenating the embeddings obtained from each of them.
4.4 The overall algorithm
Algorithm 1 outlines the whole MPS-GNN procedure for the single meta-path case (in practice, a $K$ beam size is used and multiple meta-paths are learned). The algorithm takes as input a graph ${\cal G}$, the set of available relations ${\cal R}$, an initial set of binary node labels ${\cal Y}$ and a maximal meta-path length $L_{MAX}$. It initializes the targets ${\cal S}$ with a set of singletons (one per labelled node) and their alpha values (collectively indicated by $\mathcal{A}$) to 1. At each iteration, the scoring function identifies the relation minimizing Eq. 3. The algorithm then constructs an MPS-GNN with the current (partial) meta-path, trains it on node labels, and evaluates it using the $F_{1}$ score (computed on a validation set, omitted for brevity), which reflects the meta-path’s performance when embedded in an MPS-GNN, as opposed to its potential informativeness measured by the scoring function.
Algorithm 1 MPS-GNN Learning
procedure MPS-GNN(${\cal G},{\cal R},{\cal Y},L_{MAX}$)
     $mp^{*}\leftarrow[\,]$, $F_{1}^{*}\leftarrow 0$, ${\cal S}\leftarrow{\cal Y}$, $\mathcal{A}\leftarrow 1$
     while $|mp|<L_{MAX}$ do
         $r^{*}\leftarrow\textsc{score}({\cal G},{\cal S},\mathcal{R},\mathcal{A})$
         $mp\leftarrow mp,r^{*}$
[…]
         if $F_{1}>F_{1}^{*}$ then
              $mp^{*}\leftarrow mp,\;F_{1}^{*}\leftarrow F_{1}$
         end if
         $\mathcal{A},{\cal S}\leftarrow\textsc{new-targets}({\cal S},r^{*})$
     end while
     return $mp^{*}$
end procedure
The algorithm keeps track of the best meta-path prefix together to its $F_{1}$ score, and updates the bags and $\alpha$ values as specified in Equations 5 and 6 for the next relation scoring round. The algorithm ends when the maximum meta-path length $L_{MAX}$ is reached. MPS-GNN scales linearly in the number of relations and nodes, thanks to its incremental construction of meta-paths.
4.5 MPS-GNN is a self-explainable model
By relying on meta-paths for its predictions, MPS-GNN is a self-explainable model. The scoring function serves as the detector, identifying relevant meta-paths, while the network built using them acts as the classifier. By construction, the network can only access the meta-path induced subgraph, making it strictly sufficient by construction (no changes outside the meta-path induced graph affect the prediction). An analysis of the faithfulness of MPS-GNN’s explanations is provided in Section 5.3.
5 Experiments
Our experimental evaluation seeks to address the following research questions:
  • Q1
    Can MPS-GNN recover the correct meta-path when increasing the setting complexity?
  • Q2
    Does MPS-GNN outperform existing approaches in tasks over real world relational databases?
  • Q3
    Is MPS-GNN self-explainable?
We compared MPS-GNN with approaches that don’t require predefined meta-paths, handle numerous relations, and incorporate node features in learning. The identified competitors include: MLP, to test the sufficiency of target node features alone; GCN [43], a baseline non-relational model; RGCN [44], extending GCN for multi-relational graphs, with distinct parameters for each edge type; HGN [45], a heterogeneous GNN model extending GAT for multiple relations; GTN [46], which transforms input graphs into different meta-path graphs where node representations are learned; Fast-GTN [47], an optimized GTN variant; R-HGNN [48], a relation-aware GNN using cross-relation message passing; and MP-GNN [29], the original meta-path GNN supporting only existentially quantified meta-paths.
We implemented our model using PyTorch Geometric, and used the competitors’ code from their respective papers for comparison. For training MPS-GNN, we used a 70/20/10 split for training, validation, and testing, respectively, and reported the test results for the model selected based on its validation performance. For the sake of comparison with [29], we set the maximum meta-path length to 4 and the beam size to 3. We employed $F_{1}$ as evaluation metric to account for the unbalancing in many of the datasets. The code is freely available at  [^1] https://github.com/francescoferrini/MPS-GNN
5.1 Q1: MPS-GNN consistently identifies the correct meta-path in synthetic scenarios
In order to address the first research question, we designed a sequence of synthetic node classification scenarios where the correct structure to be learnt is known. In each scenario, a node is labelled as positive if it is the starting point of at least $c$ occurrences of a given meta-path of length $l$, and negative otherwise. Crucially, existential quantification of meta-paths (as modelled by MP-GNN [29]) is insufficient here, as nodes which are starting points of less than $c$ meta-path occurrences are labelled as negatives. We designed scenarios of increasing complexity by changing the length of the ground-truth meta-path $l$, the number of occurrences $c$, and the overall number of relations $r$ in the dataset. See Figure 3 for the statistics of the different scenarios (left), and for a prototypical example for $l=2$ and $c=3$ (right).
[…]
$|\mathcal{R}|$ $c$ $l$
S1 5 2 2
S2 5 2 3
S3 5 3 3
S4 5 4 2
S5 10 2 2
S6 10 2 3
S7 10 3 3
S8 10 4 2
Table 1 shows the $F_{1}$ score of each model for an increasing complexity of the classification scenario. Results clearly show that this experimental setting is challenging for existing solutions. While the poor performance of MLP, which completely ignores the topological structure, and GCN, which ignores the difference between relations, are expected, solutions specifically conceived for heterogeneous networks also struggle with these datasets. Models like R-HGNN, HGN, GTN, and Fast-GTN, despite accounting for different relations in the graph, are affected by both the imbalance between positive and negative target nodes and the limited number of instances of neighbors of a certain type. RGCN and MP-GNN achieve better performance but are still sub-optimal. The former, like other relational methods, takes into account the diversity of relations in the graph but still uses all of them, thus struggling to single out the relevant portion of the graph. MP-GNN, on the other hand, suffers from its existential quantification assumption, and fails to find the correct meta-path in all scenarios. Conversely, MPS-GNN manages to achieve nearly-optimal performance in all scenarios, substantially outperforming all existing strategies [^2] The residual error for MPS-GNN is due to the fact that despite relying on the correct meta-path, it occasionally leverages spurious instances where the relation sequence is correct but (some of) the node features are not.. These results allow us to answer the first research question in the affirmative.
Table 1: $F_{1}$ scores on the synthetic datasets. Best results are boldfaced. Results are averaged over 5 runs. The full table including standard deviations can be found in Appendix A.4
S1 S2 S3 S4 S5 S6 S7 S8
MLP 0.46 0.44 0.48 0.47 0.44 0.51 0.45 0.47
GCN 0.46 0.46 0.48 0.52 0.44 0.48 0.46 0.48
RGCN 0.78 0.87 0.86 0.81 0.86 0.77 0.91 0.79
R-HGNN 0.50 0.44 0.47 0.47 0.53 0.48 0.46 0.48
HGN 0.45 0.46 0.50 0.46 0.46 0.48 0.45 0.40
GTN 0.46 0.52 0.49 0.48 0.44 0.47 0.49 0.47
Fast-GTN 0.46 0.48 0.51 0.49 0.44 0.47 0.53 0.47
MP-GNN 0.84 0.82 0.85 0.95 0.89 0.79 0.84 0.71
MPS-GNN 0.98 0.98 0.99 0.98 0.99 0.93 0.94 0.95
5.2 Q2: MPS-GNN surpasses competitors in real world databases, learning relevant meta-paths
Our approach is particularly useful for predictive tasks in relational databases with multiple tables, where features for a target entity may involve statistics from related tables. To address the second research question, we thus focused on three relational databases with many tables: EICU, a medical database with 31 tables, where we predict patient stay duration in the eICU; MONDIAL, a geographic database where the task is predicting whether a country’s religion is Christian; and ErgastF1, containing Formula 1 data, where the task is predicting the winner of a race. The databases were transformed into graphs as explained in Section 3; for disconnected components, we enhanced connectivity by clustering rows of auxiliary tables. Additional details for the dataset and the procedure are in the Appendix.
Results
Table 2 presents the $F_{1}$ scores of MPS-GNN and its competitors across three real-world databases, averaged over 5 runs with different seeds. The poor performance of MLP clearly indicates that using target node features only is insufficient for classification. Plain GCN, which treats the graph as homogeneous, performs well only on the EICU dataset, where node degree differences exist between positive and negative nodes. Heterogeneous GNN methods also struggle with these datasets, especially MONDIAL, where most approaches fail to outperform a simple MLP, and only one (HGN) manages to substantially outperform the non-heterogeneous baseline (GCN).
EICU MONDIAL ErgastF1
MLP 0.53 0.52 0.50
GCN 0.89 0.60 0.50
RGCN 0.70 0.53 0.57
R-HGNN 0.61 0.61 0.72
HGN 0.75 0.72 0.70
GTN 0.56 0.38 0.60
Fast-GTN 0.46 0.39 0.60
MP-GNN 0.87 0.36 0.71
MPS-GNN 0.92 0.74 0.83
Table 2: $F_{1}$-score of our method and competitors on the three real-world datasets. Full table with standard deviations can be found in Appendix A.6
MP-GNN does not provide the performance boost that was observed when applied to knowledge graphs [29], confirming our intuition that existential quantification of meta-path is insufficient when dealing with relational databases. On the other hand, MPS-GNN manages to substantially outperform all competitors, thanks to its ability to identify meta-paths that are informative thanks to the statistics that can be computed over their realizations, as shown in the following. It is worth noting that this result is achieved with one/two orders of magnitude fewer parameters than the runner-ups, namely HGN and R-HGNN. See Table 11 in the Appendix for the details. Additionally, Table 12 in the Appendix shows that MPS-GNN has competitive execution times with respect to other heterogeneous GNN approaches, thanks to its ability to focus training on relevant meta-path induced subgraphs.
Identified Meta-Paths
Figure 4: Extracted meta-paths for the three real world datasets.
Figure 4 shows the meta-paths extracted by MPS-GNN in the three real world datasets, which clearly convey relevant features for the respective task. For EICU (left), meta-paths correlate the patient’s length of stay (predictive task) with information on patients with similar periodic (top) and aperiodic (bottom) vital signs. For MONDIAL (middle), Christianity is predicted collecting information about the language of border countries (top), and the ethnic group of the country and its neighbouring countries. Finally, in ErgastF1 the winner is predicted via meta-paths collecting information about the constructor (top) and driver (top) standings.
Finally, in Appendix A.11 we present an experimental evaluation where MPS-GNN is adapted to deal with temporal databases and tasks [31], showing how it outperforms its competitors also in this context. Summing up, these results enable us to confidently answer Q2 in the affirmative.
5.3 Q3: MPS-GNN is a self-explainable method
To address the third research question, we assessed the faithfulness of the extracted meta-paths. The meta-paths identified by the model from the graph are evaluated based on the complementary metrics of sufficiency and necessity. High sufficiency implies that changing the complement to the explanation (leaving the explanation unchanged) should not affect the model’s output. High necessity implies that altering the explanation itself (leaving the complement unchanged) should result in a change in the model’s output. It is easy to show that our approach is inherently sufficient. Indeed, the computational graph of MPS-GNN consists solely of the subgraph containing the occurrences of the identified meta-paths. Necessity, on the other hand, should be experimentally evaluated by altering the computational graph of MPS-GNN.
Table 3: Changes to $F_{1}$ and necessity when removing 25%, 50%, and 75% of the learned meta-path occurrences for the real-world datasets.
$\mathbf{F_{1}}$ Necessity
Removed (%) 0 25 50 75 0 25 50 75
EICU 0.92 0.82 0.67 0.56 0 0.16 0.18 0.24
MONDIAL 0.74 0.61 0.48 0.43 0 0.13 0.29 0.32
ErgastF1 0.814 0.65 0.61 0.58 0 0.08 0.16 0.26
Table 3 illustrates the effects of removing 25%, 50%, and 75% of the meta-path occurrences in terms of changes in $F_{1}$ and necessity between original and modified graphs [49]. For a visual representation of this table and for the definition of necessity, see Appendix A.7.
In all datasets, there’s a noticeable decline in $F_{1}$ performance and a steep increase in probability difference, suggesting that the learned meta-paths are also necessary. These results clearly indicate the faithfulness of the explanations of our self-explainable method.
6 Conclusion
We introduced a novel approach to identify relevant meta-paths of relations for node classification tasks in heterogeneous graphs with a potentially large number of different relations, notably graphs derived from relational databases. Compared to earlier work, our approach does not require any user supervision, and it can learn meta-paths for predictive features defined by aggregate statistics over meta-path occurrences. Experiments clearly demonstrate advantages over alternative approaches in terms of accuracy, sparsity, and explainability. The approach was presented for binary classification, but it can be straightforwardly extended to multiclass classification using standard one-vs-all strategies. Extending it to regression tasks is an interesting direction for further research.
Appendix A Appendix
A.1 Comparison with MP-GNN
The novelty of our approach compared to MP-GNN lies in our model’s ability to learn meta-paths that are relevant to the target node class, not merely based on their existence but on statistical measures related to their occurrences. As illustrated in the example in Figure 1, MP-GNN would not be able to distinguish between the classes of the two patient nodes, as both have a single occurrence of the correct meta-path. This is due to how the function that iteratively constructs meta-paths in MP-GNN works. Specifically, under the existential quantification assumption, a candidate relation $r$ is informative for the label of a node $v$ f at least one of the neighbors ${\cal N}_{v}^{r}$ of $v$ according to $r$ belongs to the ground-truth meta-path, and $v$ has the right features. To achieve this, MP-GNN selects the optimal relation by identifying the connected node that maximizes information for class label prediction. This represents a significant limitation for MP-GNN, as it cannot differentiate between node classes that are determined by aggregate statistics (e.g. counts) of multiple occurrences of meta-paths.
A.2 Second iteration loss computation
In this Section we show the loss computation and minimization for relation "c" of the example in Figure 2. Note that $Z$ refers to the learned parameters related to the features of the target nodes in the previous iteration as shown in Figure 2 and $\binom{1}{0}$ and $\binom{0}{1}$ represent the features of $prescription$ nodes.
$Z$ $\gg 0$
$F(\mathcal{B}_{2},c)$ $=Z\Theta\binom{0}{1}+2Z\Theta\binom{1}{0}(w_{8}+w_{9})+Z\Theta\binom{1}{0}(w_{10}+w_{11})$
$F(\mathcal{B}_{3},c)$ $=2Z\Theta\binom{1}{0}(w_{8}+w_{9})+Z\Theta\binom{0}{1}(w_{10}+w_{11})+Z\Theta\binom{1}{0}(w_{12})$
$L(c)$ $=\min_{\Theta,w}\sigma\left(F(\mathcal{B}_{3},c)-F(\mathcal{B}_{2},c)\right)$
$=\min_{\Theta,w}\sigma\left(Z\Theta\binom{1}{0}(2w_{8}+2w_{9}+w_{12}-2w_{8}-2w_{9}-w_{10}-w_{11})\right.$
$\quad\left.+Z\Theta\binom{0}{1}(w_{10}+w_{11}-1)\right)$
$=\min_{\Theta,w}\sigma\left(Z\Theta\binom{1}{0}(w_{12}-w_{10}-w_{11})+Z\Theta\binom{0}{1}(w_{10}+w_{11}-1)\right)$
$\Theta\binom{1}{0}$ $\gg 0;\quad\Theta\binom{0}{1}=0;\quad w_{12}=0;\quad w_{10},w_{11}=1$
A.3 Ablation study
In this section, we highlight the importance of the skip connection $W_{1}^{(l)}h_{v}^{(0)}$ considered in equation 7. Table 4 and 5 demonstrate respectively the significance of skip connections in the synthetic and real-world settings. Without considering the initial target node features, the $F_{1}$ score drops significantly, underscoring the critical role these features play.
S1 S2 S3 S4 S5 S6 S7 S8
MPS-GNN $\textbf{0.98}\scriptstyle(\pm 0.00)$ $\textbf{0.98}\scriptstyle(\pm 0.01)$ $\textbf{0.99}\scriptstyle(\pm 0.10)$ $\textbf{0.98}\scriptstyle(\pm 0.00)$ $\textbf{0.99}\scriptstyle(\pm 0.00)$ $\textbf{0.93}\scriptstyle(\pm 0.10)$ $\textbf{0.94}\scriptstyle(\pm 0.00)$ $\textbf{0.95}\scriptstyle(\pm 0.00)$
[…]
Table 4: $F_{1}$ metric with standard deviations for synthetic datasets with MPS-GNN and MPS-GNN without using skip connection.
EICU MONDIAL ErgastF1
MPS-GNN $\textbf{0.92}\scriptstyle(\pm 0.01)$ $\textbf{0.74}\scriptstyle(\pm 0.01)$ $\textbf{0.83}\scriptstyle(\pm 0.02)$
[…]
Table 5: $F_{1}$ metric with standard deviations for real-world datasets with MPS-GNN and MPS-GNN without using skip connection.
A.4 Synthetic datasets
In table 6 are presented $F_{1}$-score with standard deviations of MPS-GNN and competitors in the synthetic settings.
S1 S2 S3 S4 S5 S6 S7 S8
MLP 0.46(±${\scriptstyle(\pm 0.00)}$ 0.44(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$ 0.47(±${\scriptstyle(\pm 0.00)}$ 0.44(±${\scriptstyle(\pm 0.00)}$ 0.51(±${\scriptstyle(\pm 0.00)}$ 0.45(±${\scriptstyle(\pm 0.00)}$ 0.47(±${\scriptstyle(\pm 0.00)}$
GCN 0.46(±${\scriptstyle(\pm 0.00)}$ 0.46(±${\scriptstyle(\pm 0.02)}$ 0.48(±${\scriptstyle(\pm 0.03)}$ 0.52(±${\scriptstyle(\pm 0.05)}$ 0.44(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$ 0.46(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$
RGCN 0.78(±${\scriptstyle(\pm 0.02)}$ 0.87(±${\scriptstyle(\pm 0.03)}$ 0.86(±${\scriptstyle(\pm 0.03)}$ 0.81(±${\scriptstyle(\pm 0.02)}$ 0.86(±${\scriptstyle(\pm 0.03)}$ 0.77(±${\scriptstyle(\pm 0.01)}$ 0.91(±${\scriptstyle(\pm 0.00)}$ 0.79(±${\scriptstyle(\pm 0.01)}$
R-HGNN 0.50(±${\scriptstyle(\pm 0.00)}$ 0.44(±${\scriptstyle(\pm 0.03)}$ 0.47(±${\scriptstyle(\pm 0.01)}$ 0.47(±${\scriptstyle(\pm 0.04)}$ 0.53(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.01)}$ 0.46(±${\scriptstyle(\pm 0.02)}$ 0.48(±${\scriptstyle(\pm 0.02)}$
HGN 0.45(±${\scriptstyle(\pm 0.00)}$ 0.46(±${\scriptstyle(\pm 0.00)}$ 0.50(±${\scriptstyle(\pm 0.03)}$ 0.46(±${\scriptstyle(\pm 0.00)}$ 0.46(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$ 0.45(±${\scriptstyle(\pm 0.03)}$ 0.40(±${\scriptstyle(\pm 0.12)}$
GTN 0.46(±${\scriptstyle(\pm 0.00)}$ 0.52(±${\scriptstyle(\pm 0.00)}$ 0.49(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$ 0.44(±${\scriptstyle(\pm 0.00)}$ 0.47(±${\scriptstyle(\pm 0.00)}$ 0.49(±${\scriptstyle(\pm 0.00)}$ 0.47(±${\scriptstyle(\pm 0.00)}$
Fast-GTN 0.46(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$ 0.51(±${\scriptstyle(\pm 0.00)}$ 0.49(±${\scriptstyle(\pm 0.00)}$ 0.44(±${\scriptstyle(\pm 0.00)}$ 0.46(±${\scriptstyle(\pm 0.00)}$ 0.53(±${\scriptstyle(\pm 0.00)}$ 0.47(±${\scriptstyle(\pm 0.00)}$
MPGNN 0.84(±${\scriptstyle(\pm 0.09)}$ 0.82(±${\scriptstyle(\pm 0.13)}$ 0.85(±${\scriptstyle(\pm 0.10)}$ 0.95(±${\scriptstyle(\pm 0.02)}$ 0.89(±${\scriptstyle(\pm 0.06)}$ 0.79(±${\scriptstyle(\pm 0.03)}$ 0.84(±${\scriptstyle(\pm 0.06)}$ 0.71(±${\scriptstyle(\pm 0.21)}$
MPS-GNN 0.98(±${\scriptstyle(\pm 0.00)}$ 0.98(±${\scriptstyle(\pm 0.01)}$ 0.99(±${\scriptstyle(\pm 0.10)}$ 0.98(±${\scriptstyle(\pm 0.00)}$ 0.99(±${\scriptstyle(\pm 0.00)}$ 0.93(±${\scriptstyle(\pm 0.10)}$ 0.94(±${\scriptstyle(\pm 0.00)}$ 0.95(±${\scriptstyle(\pm 0.00)}$
Table 6: $F_{1}$ metric with standard deviations for synthetic datasets
A.5 Real world setting
In our real-world scenario, we utilized three relational databases, which are detailed below. To convert these databases into heterogeneous graphs, we applied transformations to the attribute columns: categorical attributes were transformed using one-hot encoding, and numerical attributes were normalized to the range $[0,1]$.
To improve connectivity between target nodes, particularly when the transformation from a relational database to a graph results in each target node (or row in the target table) becoming a separate connected component, we employed simple clustering techniques on the rows of other tables based on their features.
Below, we provide detailed information about the databases and describe the clustering methods used when applicable.
EICU
Medical database with 31 tables (node types) [^3] https://eicu-crd.mit.edu. The task is predicting the duration of a patient’s stay in the eICU after admission, modeled as binary node classification by thresholding duration at 20 hours to achieve two balanced classes. To create clusters of nodes, where each cluster is represented by a single new node that replaces all the nodes within that cluster, we utilized a categorical attribute for each table that is best suited for clustering the specific table.
Table 8 provides details about the clustering process applied to the nodes of the EICU database. For each table, the initial number of rows and the resulting number of clusters (representing the final number of nodes for that type) are shown. The "Clustering Feature" column specifies the column used for creating clusters; if not specified, this indicates the absence of categorical features, and the DBSCAN algorithm is used instead.
MONDIAL
Database  [^4] https://relational-data.org/dataset/Mondial containing data from multiple geographical web data sources [50]. We predict the religion of a country as Christian (positive) with 114 instances vs. all other religions with 71 instances. In this dataset, clustering of tables is done using DBSCAN [51] clustering algorithm.
Table 9 shows the resulting number of clusters for each table of the original database. Clustering is computed using DBSCAN algorithm.
ErgastF1
Database  [^5] https://relational-data.org/dataset/ErgastF1 containing Formula 1 races from the 1950 season to the present day. It contains detailed information including lap times, pit stop durations, and qualifying performance for all races up to 2017. The objective is to predict the winner of a race using the data available before the race starts, such as the list of participants and qualifying times, while the actual lap times during the race are not yet available.
Database $|\mathcal{T}|$ $|\mathcal{R}|$ $Rows$
EICU 31 87 457325320
MONDIAL 40 45 21497
ErgastF1 14 33 544056
Table 7: Setting of real-world datasets. $|\mathcal{T}|$ and $|\mathcal{R}|$ refers respectively to the total number of tables in the original database and the number of relations in the graph used by the models. $Rows$ is the sum of all the rows of each specific database.
Table name Clustering Feature $Rows$ $Clusters$
admissiondrug drughiclseqno 7417 578
admissiondx admitdxname 7578 268
allergy drughiclseqno 2475 251
apacheapsvar - 2205 200
apachepatientresult - 3676 200
apachepredvar - 2205 200
careplancareprovider specialty 5627 40
careplaneol specialty 5627 40
careplangeneral cplgroup 3314 28
careplangoal cplcategory 3633 9
careplaninfectiousdisease cplcategory 112 11
customlab labothername 30 19
diagnosis diagnosisstring 24978 110
hospital region 186 4
infusiondrug drugname 38256 257
intakeoutput celllabel 100466 740
lab labname 434660 147
medication drughiclseqno 75604 1027
microlab organism 342 16
note notepath 24758 360
nurseassessment cellattributepath 91589 81
nursecare cellattributepath 42080 19
nursecharting nursingchartcelltypevalname 1477163 49
pasthistory pasthistorypath 12109 190
physicalexam physicalexampath 84058 310
respiratorycare currenthistoryseqnum 5436 243
respiratorycharting respchartvaluelabel 5436 243
treatment treatmentstring 38290 414
vitalaperiodic - 274088 200
vitalperiodic - 1634960 200
Table 8: Tables from the EICU dataset. ’Clustering Feature’ refers to the feature used to group the rows in each table. If not present, it means that the specific table does not have any feature for that purpose, so the DBSCAN algorithm is employed. $Clusters$ indicates the final number of nodes after the clustering step.
Table name $Rows$ $Clusters$
economy 234 5
ethnicGroup 540 65
language 144 20
politics 238 25
population 238 4
encompasses 242 2
province 1450 18
organization 153 15
continent 5 5
city 3111 93
river 218 24
sea 34 17
desert 63 6
lake 130 16
mountain 241 40
Table 9: Tables from the MONDIAL dataset. $Clusters$ indicates the final number of clusters after applying DBSCAN algorithm on the features of the specific table.
A.6 Real world setting results
EICU MONDIAL ErgastF1 rel-f1-dnf rel-f1-top3
MLP 0.53(±${\scriptstyle(\pm 0.02)}$ 0.52(±${\scriptstyle(\pm 0.00)}$ 0.50(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$
GCN 0.89(±${\scriptstyle(\pm 0.00)}$ 0.60(±${\scriptstyle(\pm 0.02)}$ 0.50(±${\scriptstyle(\pm 0.01)}$ 0.57(±${\scriptstyle(\pm 0.01)}$ 0.52(±${\scriptstyle(\pm 0.02)}$
RGCN 0.70(±${\scriptstyle(\pm 0.00)}$ 0.53(±${\scriptstyle(\pm 0.08)}$ 0.57(±${\scriptstyle(\pm 0.01)}$ 0.44(±${\scriptstyle(\pm 0.02)}$ 0.54(±${\scriptstyle(\pm 0.02)}$
R-HGNN 0.61(±${\scriptstyle(\pm 0.00)}$ 0.61(±${\scriptstyle(\pm 0.01)}$ 0.72(±${\scriptstyle(\pm 0.02)}$ 0.60(±${\scriptstyle(\pm 0.01)}$ 0.63(±${\scriptstyle(\pm 0.02)}$
HGN 0.75(±${\scriptstyle(\pm 0.00)}$ 0.72(±${\scriptstyle(\pm 0.01)}$ 0.70(±${\scriptstyle(\pm 0.04)}$ 0.61(±${\scriptstyle(\pm 0.02)}$ 0.61(±${\scriptstyle(\pm 0.01)}$
GTN 0.56(±${\scriptstyle(\pm 0.02)}$ 0.38(±${\scriptstyle(\pm 0.01)}$ 0.60(±${\scriptstyle(\pm 0.01)}$ 0.41(±${\scriptstyle(\pm 0.02)}$ 0.45(±${\scriptstyle(\pm 0.01)}$
Fast-GTN 0.46(±${\scriptstyle(\pm 0.03)}$ 0.39(±${\scriptstyle(\pm 0.04)}$ 0.60(±${\scriptstyle(\pm 0.03)}$ 0.51(±${\scriptstyle(\pm 0.01)}$ 0.50(±${\scriptstyle(\pm 0.02)}$
MP-GNN 0.87(±${\scriptstyle(\pm 0.02)}$ 0.36(±${\scriptstyle(\pm 0.06)}$ 0.71(±${\scriptstyle(\pm 0.01)}$ 0.54(±${\scriptstyle(\pm 0.02)}$ 0.52(±${\scriptstyle(\pm 0.02)}$
MPS-GNN 0.92(±${\scriptstyle(\pm 0.01)}$ 0.74(±${\scriptstyle(\pm 0.01)}$ 0.83(±${\scriptstyle(\pm 0.02)}$ 0.62(±${\scriptstyle(\pm 0.02)}$ 0.65(±${\scriptstyle(\pm 0.01)}$
Table 10: $F_{1}$-score with standard deviations of our method and competitors on real-world datasets.
A.7 Necessity calculation
To assess whether an explanation is faithful to the model, it is crucial to identify the elements that significantly influence the model’s performance. In other words, if any part of this explanation is removed, the prediction should deteriorate. This concept is known as Necessity, as it evaluates which components of the explanation are indispensable for accurate predictions. To quantify this, a distance metric is used, measuring the difference in predicted probabilities between the original predictions and those obtained after masking parts of the explanation (i.e. deleting some instances of meta-paths). This metric is calculated as:
$\textsc{Nec}=\frac{1}{N}{\sum}_{v=1}^{N}(p_{v}(\mathcal{G})-p_{v}(\mathcal{G}^{\prime}))$ (8)
where $v$ is a target node, $\mathcal{G}$ is the original graph and $p_{v}$ denotes the probability associated with the predicted class. $\mathcal{G^{\prime}}$ is obtained by removing certain meta-path occurrences and $p_{v}(\mathcal{G^{\prime}})$ is the probability associated at the class predicted with $p_{v}$. Since MPS-GNN utilizes just the explanation for making prediction one should expect that removing some part of the explanation, has as effect a decrease in prediction $F_{1}$. Figure 5 shows five plots, one for each dataset. As in Table 3 each graph illustrates the effects of removing 25%, 50%, and 75% of the meta-path occurrences in terms of changes in $F_{1}$ and necessity between original and modified graphs [49].
Figure 5: Changes to $F_{1}$ and posterior probability difference (necessity) when removing 25%, 50%, and 75% of the learned meta-path occurrences for the real-world datasets.
A.8 Number of parameters
In Table 11, we present the total number of parameters required for evaluating the various models. In the synthetic setting, when comparing with the only two models that yield satisfactory results, we observe that our approach has a similar number of parameters as RGCN (when the total number of relations in the graphs is limited) and MP-GNN. MP-GNN, which also considers only a subset of graph relations like our method, is designed to have a lower number of parameters; however, it still falls short of matching MPS-GNN’s performance.
In the real-world setting, among the models that achieve decent results, GCN exhibits the lowest number of parameters on the EICU dataset. However, among the relational methods, MPS-GNN emerges as the most efficient. On the MONDIAL dataset, the two leading competitors, HGN and R-HGNN, despite achieving lower $F_{1}$ scores, utilize all edge types and consequently require a significantly larger number of parameters.
Finally, on the ErgastF1 dataset, although MP-GNN outperforms other methods in terms of parameter efficiency by considering different meta-paths, it results in a considerably lower $F_{1}$ score. In contrast, HGN and R-HGNN exhibit an exponential increase in the number of parameters.
MLP GCN RGCN R-HGNN HGN GTN Fast-GTN MP-GNN MPS-GNN
S1 194 690 3730 525996 10927 866 126902 1346 3618
S2 194 690 6770 787796 42314 946 126942 1346 3618
S3 194 690 3730 525996 42314 866 126902 1346 3618
S4 194 690 6770 787796 74125 946 126942 1346 3618
S5 194 690 3730 525996 74125 866 126902 1346 6786
S6 194 690 6770 787796 74125 946 126942 1346 6786
S7 194 690 3730 525996 74125 866 126902 1346 6786
S8 194 690 6770 787796 74125 946 126942 8834 6786
EICU 1346 3506 400690 47496672 611785 32024 110408 24898 12066
MONDIAL 2144 90546 3709106 88396572 1142962 180554 1539457 183138 234050
ErgastF1 4356 198142 5422432 396543021 439021 2542354 11325242 23413 29538
Table 11: Number of parameters for each model across synthetic and real-world datasets.
A.9 Execution times
In Table 12, we present the training times for each model across the individual datasets. In the synthetic settings, we observe that among models achieving a significant $F_{1}$ score (RGCN, MP-GNN, and MPS-GNN), MPS-GNN typically demonstrates the shortest execution time. We would like to highlight that our model is specifically designed to learn meaningful meta-paths in networks with many relation types, whereas the synthetic datasets are limited in the number of relation types. In the EICU database, while the MLP model achieves the shortest execution time, it performs poorly in terms of $F_{1}$. Among the models with notable results, MPS-GNN exhibits the best execution time. In the MONDIAL database, all models have relatively low execution times due to the small graph size, as shown in Table 7. However, MPS-GNN still achieves the best $F_{1}$ score. For the ErgastF1 dataset, while MLP again has the lowest execution time, its final accuracy is poor. In contrast, MPS-GNN is comparable to MP-GNN and R-HGNN in terms of execution time but surpasses them by $12$ and $11$ percentage points in $F_{1}$ score, respectively. Overall, our approach is consistently neither the quickest nor the slowest, yet it reliably achieves the highest average F1 score across all settings.
MLP GCN RGCN R-HGNN HGN GTN Fast-GTN MP-GNN MPS-GNN
S1 66 660 1094 898 363 388 315 234 322
S2 67 660 2197 674 375 570 812 1461 245
S3 73 675 536 612 380 260 456 657 356
S4 72 483 2142 551 373 697 369 986 457
S5 69 420 445 575 360 875 845 158 467
S6 69 620 111 616 377 567 467 453 321
S7 72 677 285 519 371 834 442 587 449
S8 74 777 167 680 373 878 765 1502 490
EICU 342 5882 4355 4842 3210 10324 682 5787 1273
MONDIAL 156 125 132 131 265 220 119 120 134
ErgastF1 543 954 1491 2456 2245 3015 2945 2280 2421
Table 12: Training times, in seconds, for each model across synthetic and real-world datasets.
A.10 Scoring Function Lookahead Illustration
We report an additional experiment demonstrating the lookahead capabilities implemented in the scoring function of our method, and its advantage over a simplistic greedy approach. We construct a synthetic dataset of a multi-relational graph with three relations and the ground truth metapath, $r_{1},r_{2}$ for the target. We compare our approach against a simple greedy one, in which one always extends the metapath with the relation for which a trained MPS-GNN achieves maximal $F_{1}$ score. The results are presented in Table 13. In the first iteration of the scoring function, relation $r_{1}$ achieves the lowest loss in our scoring function and would therefore be chosen as the first relation of the metapath. Looking only at the immediate benefit of the relations in terms of the accuracy achieved by a corresponding MPS-GNN, however, $r_{2}$ would be selected as the best relation. The column $r_{2}$-Extensions shows the $F_{1}$ scores of all possible length 2 metapaths starting with $r_{2}$. Comparing with the $F_{1}$ score of the ground truth metapath we find that, indeed, starting the metapath with $r_{2}$ is a suboptimal choice, and that our scoring function correctly identifies the most informative relation to start the metapath with, even though this informativeness only is materialized after extension of the metapath with $r_{2}$.
Iteration 1 $r_{2}$- Extensions Ground truth
Meta-paths $r_{1}$ $r_{2}$ $r_{3}$ $r_{2},r_{1}$ $r_{2},r_{2}$ $r_{2},r_{3}$ $r_{1},r_{2}$
Score 0.001 45 56
$F_{1}$ 0.79 0.82 0.69 0.83 0.82 0.85 0.99
Table 13: Comparison: our metapath construction vs. simple greedy alternative.
A.11 Temporal experiment
Recently, a novel benchmark, rel-bench [31], has been introduced. This benchmark consists of multiple relational databases represented as temporal graphs. Additionally, the authors propose RDL, a temporal-aware relational message-passing model. It’s important to note that the released temporal graphs are not directly compatible with static models. Therefore, in this section, we reconstruct one dataset from the rel-bench repository in a static format and apply our method to this static representation of the graph (setting in table 14). The tasks are: (1) dnf, predicting whether a driver will fail to finish a race within the next month, and (2) top3, predicting if a driver will place in the top 3. To handle these temporal tasks, we treated each instance of a node at different timestamps as distinct nodes with separate label predictions. Table 15 reports the $F_{1}$-scores for MPS-GNN and the baselines. Our approach outperforms all competitors, including RDL which performs lower due to overpredicting the majority class. Note that RDL cannot be straightforwardly applied to our other datasets, being designed for temporal datasets. In Section A.11.1, we provide the necessity calculations for these datasets.
Database $|\mathcal{T}|$ $|\mathcal{R}|$ $Rows$
rel-f1 9 26 97606
Table 14: Setting of temporal real-world dataset.
rel-f1-dnf rel-f1-top3
MLP $0.48{\scriptstyle(\pm 0.00)}$ $0.48{\scriptstyle(\pm 0.00)}$
GCN $0.57{\scriptstyle(\pm 0.01)}$ $0.52{\scriptstyle(\pm 0.02)}$
RGCN $0.44{\scriptstyle(\pm 0.02)}$ $0.54{\scriptstyle(\pm 0.02)}$
R-HGNN $0.60{\scriptstyle(\pm 0.01)}$ $0.63{\scriptstyle(\pm 0.02)}$
HGN $0.61{\scriptstyle(\pm 0.02)}$ $0.61{\scriptstyle(\pm 0.01)}$
GTN $0.41{\scriptstyle(\pm 0.02)}$ $0.45{\scriptstyle(\pm 0.01)}$
Fast-GTN $0.51{\scriptstyle(\pm 0.01)}$ $0.50{\scriptstyle(\pm 0.02)}$
MP-GNN $0.54{\scriptstyle(\pm 0.02)}$ $0.52{\scriptstyle(\pm 0.02)}$
RDL $0.58{\scriptstyle(\pm 0.03)}$ $0.53{\scriptstyle(\pm 0.7)}$
MPS-GNN 0.62(±${\scriptstyle(\pm 0.02)}$ 0.65 (±${\scriptstyle(\pm 0.01)}$
Table 15: $F_{1}$-score with standard deviations of our method and competitors on two temporal datasets.
A.11.1 Necessity calculation in temporal tasks
In this section, we present the necessity calculation for the temporal datasets. Specifically, as detailed in Appendix A.7, Table 16 and figure6 demonstrates that removing certain identified meta-paths results in decreased performance ($F_{1}$) and an increased necessity value. This finding confirms that the learned meta-paths are essential for the prediction task in these datasets as well.
$\mathbf{F_{1}}$ Necessity
Removed (%) 0 25 50 75 0 25 50 75
rel-f1-dnf 0.63 0.56 0.49 0.48 0 0.12 0.22 0.31
rel-f1-top3 0.65 0.62 0.55 0.50 0 0.19 0.34 0.34
Table 16: Changes to $F_{1}$ and necessity when removing 25%, 50%, and 75% of the learned meta-path occurrences for the real-world temporal tasks rel-f1-dnf and rel-f1-top3.
Figure 6: Changes to $F_{1}$ and posterior probability difference (necessity) when removing 25%, 50%, and 75% of the learned meta-path occurrences for the real-world temporal tasks rel-f1-dnf and rel-f1-top3.
A.12 Non-GNN models on real-world databases
For the MONDIAL and ErgastF1 databases, non-GNN methods have shown competitive performance in the past. For example, [52, 53] report $F_{1}$ scores of $0.78$ and $0.77$ on MONDIAL, $0.4$ and $0.3$ points higher than MPS-GNN. However, these results are achieved on a simplified version of the database with only 12 tables, requiring manual feature selection. In contrast, MPS-GNN is applied directly to the raw input data across all 40 tables. The non-GNN methods use Multi-relational Bayes Net Classifiers and Simple Decision Forests, where reducing the number of tables and relations aids performance. By comparison, MPS-GNN is designed to handle scenarios with a large number of relations effectively.

### v2

@@
 1 Introduction
-Graph Neural Networks (GNNs) have increasingly become the de-facto standard for many predictive tasks involving networked data, such as physical systems [1, 2], Knowledge Graphs [3] and social networks [4].
+Graph Neural Networks (GNNs) have increasingly become the de-facto standard for many predictive tasks involving networked data, such as physical systems (Sanchez-Gonzalez et al., 2018; Battaglia et al., 2016), Knowledge Graphs (Hamaguchi et al., 2017) and social networks (Wu et al., 2020).
 By learning effective node embeddings, GNNs offer a unified framework for addressing various graph-based tasks, including node classification, graph classification, and link prediction.
 However, similar to other representation learning paradigms, GNNs often exhibit a black-box nature in their predictions.
-Numerous solutions have been proposed for post-hoc explainability of GNN predictions, primarily at the instance-based level [5, 6, 7, 8].
-Yet, as with other deep learning architectures, the ability of these approaches to genuinely reflect the underlying reasoning of the predictor has been called into question [9].
-To tackle this challenge, self-explainable GNNs [10, 11, 12] have recently emerged, aiming to ensure that GNN predictions are grounded in interpretable elements, such as subgraphs [13, 14] or prototypes [15, 16].
+Numerous solutions have been proposed for post-hoc explainability of GNN predictions, primarily at the instance-based level (Ying et al., 2019; Vu & Thai, 2020; Miao et al., 2022; Yuan et al., 2021).
+Yet, as with other deep learning architectures, the ability of these approaches to genuinely reflect the underlying reasoning of the predictor has been called into question (Longa et al., 2024).
+To tackle this challenge, self-explainable GNNs (Kakkad et al., 2023; Christiansen et al., 2023; Seo et al., 2023) have recently emerged, aiming to ensure that GNN predictions are grounded in interpretable elements, such as subgraphs (Wu et al., 2022; Yu et al., 2020) or prototypes (Zhang et al., 2022; Ragno et al., 2022).
 Despite the widespread adoption of GNNs, most approaches are tailored for homogeneous graphs, where edge types are indistinguishable.
@@
 This limitation is particularly problematic in knowledge graphs, which typically feature numerous relations (corresponding to edge types), with only a few being pertinent to a specific predictive task—forming the so-called meta-paths.
-Existing approaches for heterogeneous GNNs either rely on domain experts to provide relevant meta-paths a priori [17, 18, 19, 20], or attempt to learn them from data by assigning different weights to various relations  [21, 22, 23, 24, 25, 26, 27, 28] , a solution that fails to scale with the number of candidate relations.
-Recently, MP-GNN [29] has been proposed as a solution for learning meta-paths without requiring user supervision.
+Existing approaches for heterogeneous GNNs either rely on domain experts to provide relevant meta-paths a priori (Chang et al., 2022; Fu et al., 2020; Li et al., 2021; Wang et al., 2019), or attempt to learn them from data by assigning different weights to various relations  (Hu et al., 2020; Lv et al., 2021b; Mitra et al., 2022; Schlichtkrull et al., 2018; Yu et al., 2022; Yun et al., 2019b; 2022a; Zhu et al., 2019) , a solution that fails to scale with the number of candidate relations.
+Recently, MP-GNN (Ferrini et al., 2024) has been proposed as a solution for learning meta-paths without requiring user supervision.
 This approach employs a scoring function to predict the potential informativeness of partial meta-paths, enabling efficient exploration of the combinatorial space of candidate meta-paths.
-However, a key limitation of MP-GNN is its assumption that a node’s class is primarily dependent on the existence of a meta-path instance.
+However, a key limitation of MP-GNN is the existential quantification assumption which states that a node’s class is primarily dependent on the existence of a meta-path instance, meaning just one instance is sufficient for accurate prediction.
+its assumption that a node’s class is primarily dependent on the existence of a meta-path instance, meaning just one instance is sufficient for accurate prediction.
 While this assumption may be reasonable for knowledge graphs, it is unrealistic when dealing with relational databases, where entities are characterized by numerous categorical and numerical attributes.
-On the other hand, this same complexity makes relational databases particularly well-suited for GNN technology, as evidenced by the growing interest in what is now being termed relational deep learning [30].
+On the other hand, this same complexity makes relational databases particularly well-suited for GNN technology, as evidenced by the growing interest in what is now being termed relational deep learning (Fey et al., 2023).
 Figure 1: Left: Relational database schema for a medical domain.
 Right: Heterogeneous graph representation of (part of) the database.
-The highlighted subgraph shows a prototypical counts-of-counts pattern characterising positive patients, namely having at least two exempt prescriptions (indicated as true), each containing at least two medications.
-Existing heterogeneous GNNs struggle with these patterns as they need to learn a separate weight matrix for each edge type in the graph, while MPS-GNN is capable of learning them without any direct user supervision.
+The highlighted subgraph shows a prototypical counts-of-counts pattern characterising positive patients, namely having at least two exempt prescriptions (represented by node feature T), each containing at least two medications.
+Existing heterogeneous GNNs struggle with these patterns as they need to learn a separate weight matrix for each edge type in the graph, while MPS-GNN is capable of learning the relevant meta path without any direct user supervision.
 In this paper, we extend the concept behind MP-GNN beyond the simple existential quantification of meta-paths.
@@
 These include counts-of-counts statistics such as having at least two exempt prescriptions with at least two medications each to characterize patients with severe illness, as shown in Figure 1.
+Note that MP-GNN would fail to discriminate between the two patient nodes in the figure, as they both have at least one occurrence of the correct meta-path.
 An experimental evaluation on both synthetic and real-world relational database tasks demonstrate the significant advantages of the proposed solution over existing alternatives.
-Additionally, results show how the meta-path learning strategy behind MPS-GNN renders it inherently and genuinely self-explainable, in contrast to many existing self-explainable GNN architectures whose explanations often lack fidelity [11].
+Additionally, results show how the meta-path learning strategy behind MPS-GNN renders it inherently and genuinely self-explainable, in contrast to many existing self-explainable GNN architectures whose explanations often lack fidelity (Christiansen et al., 2023).
 2 Related Work
-Relational deep learning [30] has recently emerged as a paradigm advocating the application of deep learning technology, and GNNs in particular, to relational databases.
-The rationale behind this research direction is the popularity of relational databases as a mean to store relational information in a variety of application domains, combined with the fact that relational databases can be seen as heterogeneous graphs, with tables converted into sets of nodes and relations into (typed) edges between table entries [31].
+Relational deep learning (Fey et al., 2023) has recently emerged as a paradigm advocating the application of deep learning technology, and GNNs in particular, to relational databases.
+The rationale behind this research direction is the popularity of relational databases as a mean to store relational information in a variety of application domains, combined with the fact that relational databases can be seen as heterogeneous graphs, with tables converted into sets of nodes and relations into (typed) edges between table entries (Robinson et al., 2024).
 This transformation allows the application of heterogeneous Graph Neural Networks (GNNs) to this kind of data.
@@
 Existing approaches to incorporate meta-paths follow in two main categories.
-The former requires domain experts to identify relevant meta-paths for the task at hand [17, 18, 19, 20].
+The former requires domain experts to identify relevant meta-paths for the task at hand (Chang et al., 2022; Fu et al., 2020; Li et al., 2021; Wang et al., 2019).
 Clearly, this approach is suboptimal as it requires this domain information to be readily available.
-The alternative solutions, either utilize relation-specific graph convolutions, capturing relational patterns with distinct parameters or dedicated components for each type of relation  [21, 22, 24, 25] or focuses on graph transformation and multi-view learning to enhance relational representations [23, 26, 27, 28].
+The alternative solutions, either utilize relation-specific graph convolutions, capturing relational patterns with distinct parameters or dedicated components for each type of relation  (Hu et al., 2020; Lv et al., 2021b; Schlichtkrull et al., 2018; Yu et al., 2022) or focuses on graph transformation and multi-view learning to enhance relational representations (Mitra et al., 2022; Yun et al., 2019b; 2022a; Zhu et al., 2019).
 Although these methods are effective with a limited number of relations, their performance quickly deteriorates as the number of candidate relations grows.
-Recently, a novel approach named MP-GNN [29] has been introduced to tackle the aforementioned challenges and automatically learn relevant meta-paths from data.
+Recently, a novel approach named MP-GNN (Ferrini et al., 2024) has been introduced to tackle the aforementioned challenges and automatically learn relevant meta-paths from data.
 The approach leverages a scoring function predicting the potential informativeness of partial meta-paths to guide the search in the combinatorial space of candidate meta-paths.
@@
 This extension is crucial in allowing MPS-GNN to be effectively applied to relational deep learning settings, as shown by our experimental evaluation.
-More details about the comparison between MP-GNN and MPS-GNN in A.1.
-Explainability in GNN is a major research trend, with plenty of approaches for post-hoc explanation of GNN predictions (instance-based explainability [5, 6, 7, 8]) and others aiming to explain the GNN model as a whole (model-based explainability [32, 33, 34, 35]).
-Specialized metrics have been developed to estimate the faithfulness of an explanation in relation to the method’s input processing behavior [36, 37, 38, 39, 40].
-Following a similar evolution in the XAI literature of convolutional neural networks, self-explainable GNNs [13, 14, 15, 16] have recently been proposed to encourage GNN models to adhere to their explanations by design.
-However, experimental studies have questioned the faithfulness of the explanations provided by these approaches [11, 41], highlighting the difficulty of achieving genuine explainability with GNNs.
+More details about the comparison between MP-GNN and MPS-GNN in 4.6.
+Explainability in GNN is a major research trend, with plenty of approaches for post-hoc explanation of GNN predictions (instance-based explainability (Ying et al., 2019; Vu & Thai, 2020; Miao et al., 2022; Yuan et al., 2021)) and others aiming to explain the GNN model as a whole (model-based explainability (Chen et al., 2024; Wang & Shen, 2022; Azzolin et al., 2022; Yuan et al., 2020)).
+Specialized metrics have been developed to estimate the faithfulness of an explanation in relation to the method’s input processing behavior (Agarwal et al., 2023; Yuan et al., 2022; Amara et al., 2022; 2023; Zheng et al., 2023).
+Following a similar evolution in the XAI literature of convolutional neural networks, self-explainable GNNs (Wu et al., 2022; Yu et al., 2020; Zhang et al., 2022; Ragno et al., 2022) have recently been proposed to encourage GNN models to adhere to their explanations by design.
+However, experimental studies have questioned the faithfulness of the explanations provided by these approaches (Christiansen et al., 2023; Azzolin et al., 2025), highlighting the difficulty of achieving genuine explainability with GNNs.
 A key advantage of our proposed method is that meta-paths can naturally serve as model-level explanations, making MPS-GNN the first truly self-explainable GNN designed for relational deep learning applications.
@@
 Definition 3.1 (Relational Database).
-A relational database $(\mathcal{T},\mathcal{L})$ consists of a collection of tables $\mathcal{T}=T_{1},...T_{n}$ and links between these tables $\mathcal{L}\subseteq T\times T$ Each table is a set $T=\{u_{1},...,u_{n}\}$ where the elements $u\in T$ are referred to as rows or entities.
-Each entity is a tuple $u=(\mathcal{P}_{u},\mathcal{K}_{u},a_{u})$ where $\mathcal{P}_{u}$ is the Primary Key that uniquely identifies the entity $u$; $\mathcal{K}_{u}$ is the set of Foreign Keys corresponding to a primary key in other tables, thus connecting the tables; $a_{u}$ corresponds to the Attributes of the entity $u$.
+A relational database $(\mathcal{T},\mathcal{L})$ consists of a collection of tables $\mathcal{T}=T_{1},...T_{n}$ and links between these tables $\mathcal{L}\subseteq T\times T$ Each table is a set $T=\{e_{1},...,e_{n}\}$ where the elements $e\in T$ are referred to as rows or entities.
+Each entity is a tuple $e=(\mathcal{P}_{e},\mathcal{K}_{e},a_{e})$ where $\mathcal{P}_{e}$ is the Primary Key that uniquely identifies the entity $e$; $\mathcal{K}_{e}$ is the set of Foreign Keys corresponding to a primary key in other tables, thus connecting the tables; $a_{e}$ corresponds to the Attributes of the entity $e$.
 Definition 3.2 (Heterogeneous graph).
@@
 knowledge graphs.
-From relational database to heterogeneous graph A relational database can be interpreted as an heterogeneous graph where row $u$ becomes node $v$; attributes $a_{u}$ become node features $x_{v}$; links $\mathcal{L}$ between entries of two tables are identified by the pair of primary $\mathcal{P}$ and foreign $\mathcal{K}$ keys in the two tables.
+From relational database to heterogeneous graph A relational database can be interpreted as an heterogeneous graph where row $e$ becomes node $v$; attributes $a_{e}$ become node features $x_{v}$; links $\mathcal{L}$ between entries of two tables are identified by the pair of primary $\mathcal{P}$ and foreign $\mathcal{K}$ keys in the two tables.
 Each pair of connected tables, originates a relation in the graph, specified by $r$.
 4 Methodology
+Figure 2: Outline of greedy local meta-path construction
 For now we restrict attention to binary node classification problems.
@@
 We construct meta-paths following a strategy that is
-  • Greedy: a partially constructed meta-path $\xrightarrow{r_{1}}\xrightarrow{r_{2}}\ldots\xrightarrow{r_{i}}$ is extended by a next relation $\xrightarrow{r_{i+1}}$ without lookahead for possible completions $\xrightarrow{r_{i+1}}\ldots\xrightarrow{r_{L}}$.
-Similar as in [29], we try to estimate the potential informativeness of nodes reached by $\xrightarrow{r_{i+1}}$ by learned weights associated with the nodes.
+  • Greedy: a partially constructed meta-path $\xrightarrow{r_{1}}\xrightarrow{r_{2}}\ldots\xrightarrow{r_{i}}$ is extended by a next relation $\xrightarrow{r_{i+1}}$ without lookahead for possible completions $\xrightarrow{r_{i+2}}\ldots\xrightarrow{r_{L}}$.
+Similar as in Ferrini et al.
+(2024), we try to estimate the potential informativeness of nodes reached by $\xrightarrow{r_{i+1}}$ by learned weights associated with the nodes.
 These weights represent putative features that can either be directly materialized as functions of the nodes’ attributes, or that can be constructed as features of meta-path extensions starting at the node.
@@
 The surrogate problems take the form of weighted multi-instance classification tasks.
+Figure2 illustrates the high-level principles of our approach.
+Given an initial (binary) node classification task, a first relation is identified that could solve the task with the help of a putative node feature (weight) $\boldsymbol{w}$ on the successor nodes.
+Then a surrogate classification task is set up whose target is to materialize the putative feature as a feature computable from the data.
+This surrogate task takes the form of a weighted multi-instance classification task, which can be seen as an abstraction of a direct regression problem with target $\boldsymbol{w}$ (see Section 4.2).
 Note that both the greedy and local properties mirror core principles of growing decision trees, which are built incrementally, adding one relation at a time based on solving local classification sub-tasks.
+The improvements in time complexity with respect to an algorithm that looks at every possible relations in the meta-path construction are detailed in section 4.2.1.
+In Section 4.1, we describe the weighted multi-instance classification task and its application in our scenario.
+Section 4.2 outlines the methodology for constructing meta-paths by scoring graph relations, including the examples based on Figure 1.
+In Section 4.3, we detail the GNN used in our setup, and Section 4.4 presents the complete framework of the proposed approach.
 4.1 Weighted multi-instance classification
@@
 Our goal is to predict the bag label via discriminant functions of the form
-$F({\mathcal{B}})=\sum_{v\in{\mathcal{B}}}\alpha(v,{\mathcal{B}})\cdot F(v),$ (1)
+$F({\mathcal{B}})=\sum_{v\in{\mathcal{B}}}\alpha(v,{\mathcal{B}})f(v),$ (1)
 where $f(v)$ is a learnable node feature function. Specifically, we consider functions that are parameterized by a relation $r$, and are of the form
-$F(v,r,\Theta,{\boldsymbol{w}})=\Theta\cdot x_{v}\cdot\sum_{u\in\mathcal{N}^{r}_{v}}w_{u},$ (2)
+$f(v,r,\Theta,{\boldsymbol{w}})=\left\{\begin{array}[]{ll}\Theta^{T}x_{v}&\mbox{if}\ \mathcal{N}^{r}_{v}=\emptyset\\ \Theta^{T}x_{v}\sum_{u\in\mathcal{N}^{r}_{v}}w_{u}&\mbox{if}\ \mathcal{N}^{r}_{v}\neq\emptyset\end{array}\right.$ (2)
 where $x_{v}$ denotes the attribute vector of $v$, $\Theta$ is a trainable parameter vector, and ${\boldsymbol{w}}$ is a vector of trainable weights $w_{u}\in[0,1]$ assigned to $v$’s $r$-neighbors.
-The node feature function value for $v$ can, thus, be determined by $v$’s own attributes, or by aggregating the putative feature $w_{u}$ of its successors – or a combination of both elements.
+The node feature function is thus computed as a combination of $v$’s own attributes and the putative feature $w_{u}$ of its successors.
+When the node attributes $x_{v}$ are not informative for solving the multi-instance classification task, then parameters $\Theta$ can be learned that make $\Theta^{T}x_{v}$ constant for all $v$, and thus this part of the node feature function becomes irrelevant (a suitable $\Theta$ exists e.g.
+under the mild assumption that $x_{v}$ contains at least one categorical attribute  [^1] In a heterogeneous graph built from a relational database, this would be the table the node belongs to.
+If node representations are simply embeddings, one can achieve the same goal by concatenating a constant feature to the embedding.
+in a one-hot encoding: then $\Theta^{T}x_{v}=1$ for $\Theta$ that is set to 1 for all entries corresponding to the one-hot encoding of the categorical attribute, and zero everywhere else).
+The factor $\sum_{u\in\mathcal{N}^{r}_{v}}w_{u}$ captures a dependence of $f(v)$ on the $r$-neighborhood of $v$.
+When the $r$-neighborhood is non-informative because neither the number nor the identity of nodes $u\in\mathcal{N}^{r}_{v}$ has discriminative value, then the $\sum_{u\in\mathcal{N}^{r}_{v}}w_{u}$ factor can be made irrelevant by learning constant weights $w_{u}$.
 Ideally, the discriminant function separates the classes in the sense that for any pair ${\mathcal{B}}^{+}\in{\mathcal{S}}^{+},{\mathcal{B}}^{-}\in{\mathcal{S}}^{-}$ we have $F({\mathcal{B}}^{+})>F({\mathcal{B}}^{-})$.
@@
 $L(r,\Theta,{\boldsymbol{w}})=\sum_{{\mathcal{B}}^{+}\in\mathcal{S}^{+},{\mathcal{B}}^{-}\in\mathcal{S}^{-}}\sigma\Big{(}F({\mathcal{B}}^{-},r,\Theta,{\boldsymbol{w}})-F({\mathcal{B}}^{+},r,\Theta,{\boldsymbol{w}})\Big{)},$ (3)
-where $\sigma$ is the sigmoid function. In practice, we approximate (3) by a random sample of positive and negative bags.
+where $\sigma$ is the sigmoid function.
+In practice, given that the number of terms in the sum is quadratic in the number of training examples, we approximate (3) by a random sample of positive and negative bags.
 4.2 Relation Scoring
@@
 Denote with ${\mathcal{S}}_{1}^{+},{\mathcal{S}}_{1}^{-}$ the sets of all initial positive and negative bags, respectively.
-At all iterations we select the relation that minimizes the loss
+At all iterations we select the relation that minimizes the loss (also referred to as the scoring function)
 $L(r)=\min_{\Theta,{\boldsymbol{w}}}L(r,\Theta,{\boldsymbol{w}}).$ (4)
-If for some relation $r$ this loss is effectively minimized by suitable parameters $\Theta$, such that the $\sum w_{u}$ term in (2) plays no significant role, then the meta-path construction can terminate at this point (without adding $r$ to the meta-path).
-If, on the other hand, the $r$-successors and their weights ${\boldsymbol{w}}$ are needed to solve the current multi-instance classification task, then we need to extend the meta-path by $r$, and we need to capture the putative node features ${\boldsymbol{w}}$ by an actual node feature.
+If all candidate relations fail to minimize the loss, i.e., optimizing $\Theta,{\boldsymbol{w}}$ does not lead to substantially lower loss than using random parameters (i.e., does not improve by at least 30%), then the meta-path construction terminates and the current meta-path is returned.
+Otherwise, the current meta-path is extended with the minimal loss relation $r$.
+At this point, the problem becomes capturing the putative node features ${\boldsymbol{w}}$ by actual features.
 A possible approach would be to set this up as a node regression task with target values $w_{u}$.
@@
 For this, let $r_{i}$ denote the relation found to minimize (4) in iteration $i$ with optimal parameters $\Theta_{i}$.
-For each positive bag ${\mathcal{B}}^{+}_{i}\in\mathcal{S}_{i}^{+}$ define the new bag
-${\mathcal{B}}^{+}_{i+1}=\cup_{v\in{\mathcal{B}}^{+}_{i}}\mathcal{N}^{r_{i}}_{v}$ (5)
-containing the $r_{i}$-children of the nodes in ${\mathcal{B}}^{+}_{i}$. Similarly for negative bags. For $u\in{\mathcal{B}}^{+}_{i+1}$ define the node weight
-$\alpha(u,{\mathcal{B}}^{+}_{i+1})=\sum_{v\,:\,u\in\mathcal{N}^{r_{i}}_{v}}\Theta_{i}\cdot x_{v}\cdot\alpha(v,{\mathcal{B}}^{+}_{i}).$ (6)
-Thus, nodes have a bigger weight in the subsequent multi-instance classification task, if they are $r_{i}$- successors of many nodes from the previous task, that make a significant contribution to the discriminant function (1) learned for the previous task.
-The sets of all ${\mathcal{B}}^{+}_{i+1}({\mathcal{B}}^{-}_{i+1})$ form the new training sets ${\mathcal{S}}^{+}_{i+1}({\mathcal{S}}^{-}_{i+1})$.
+For each positive bag ${\mathcal{B}}^{+}\in\mathcal{S}_{i}^{+}$ define the new bag
+${\mathcal{B}}^{+}_{\emph{new}}=\cup_{v\in{\mathcal{B}}^{+}}\mathcal{N}^{r_{i}}_{v}$ (5)
+containing the $r_{i}$-children of the nodes in ${\mathcal{B}}^{+}$.
+Similarly for negative bags.
+For $u\in{\mathcal{B}}^{+}_{\emph{new}}$ define the node weight by the following sum over all nodes $v\in\mathcal{B}^{+}$ that have $u$ as an $r_{i}$-neighbor:
+$\alpha(u,{\mathcal{B}}^{+}_{\emph{new}})=\sum_{\{v\in\mathcal{B}^{+}\,:\,u\in\mathcal{N}^{r_{i}}_{v}\}}\Theta_{i}^{T}x_{v}\alpha(v,{\mathcal{B}}^{+}).$ (6)
+This definition of the weights enables to essentially ignore in the setup of the multi-instance classification task for iteration $i+1$ those nodes $u$ that are $r_{i}$-successors only of nodes $v$ that did not play a major role in solving the task of the previous iteration – either because of a low absolute value of $\Theta_{i}^{T}x_{v}$, or because $\alpha(v,{\mathcal{B}}^{+}_{i})$ already had a low absolute value.
+The sets of all ${\mathcal{B}}^{+}_{\emph{new}}({\mathcal{B}}^{-}_{\emph{new}})$ form the new training sets ${\mathcal{S}}^{+}_{i+1}({\mathcal{S}}^{-}_{i+1})$.
+Toy Example:
+To provide a clear understanding of the proposed approach, we illustrate it with the toy example of Figure 1, where patients are positive if, and only if, they have at least two exempt prescriptions, each containing at least two medications.
+Initially, the approach evaluates the potential informativeness of two relations originating from patients: relation $a$ and relation $b$.
+Figure 3(a) illustrates the scoring process for relation a.
+The first step involves instantiating the node feature functions for patient nodes 0 and 1 by applying Equation 2.
+Since both nodes share the same unique neighbor, the loss function 4 reduces to the constant $\frac{1}{2}$, which cannot be minimized by $\Theta$ and ${\boldsymbol{w}}$.
+(a) Scoring relation $a$
+(b) Scoring relation $b$
+Figure 3: Scoring the first two relations
+Similarly, relation $b$ is scored. The loss function 4 now becomes
+$\sigma(\Theta^{T}x_{1}(w_{4}+w_{5}+w_{6})-\Theta^{T}x_{0}(w_{2}+w_{3}+w_{4})),$
+where $x_{0}=x_{1}=(1,0)^{T}$ is the attribute vector representing the T value in a one-hot encoding.
+This loss can be brought arbitrarily close to zero, e.g.
+by $w_{6}\rightarrow 0$, $w_{5}\rightarrow 0$, $w_{3}\rightarrow 1$, $w_{2}\rightarrow 1$, and $\Theta=(Z,0)^{T}$ with $Z\gg 0$.
+Consequently, relation $b$ scores higher than relation $a$ and is selected for further extension.
+Note that the low loss of relation $b$ is entirely due to its potential informativeness: the $b$ neighborhoods of the nodes 0 and 1 are completely isomorphic, and therefore the meta path consisting of $b$ alone provides no discriminative information.
+To extend the meta path prefix $b$ we set up the weighted multi-instance classification task for the second iteration.
+This gives us a positive bag $\mathcal{B}_{2}$ containing nodes $\{2,3,4\}$ and a negative bag $\mathcal{B}_{3}$ containing nodes $\{4,5,6\}$.
+Node weights for the bags are computed according to Equation (6), and here simply yield uniform weights $\alpha(k,\mathcal{B}_{2})=Z$ for all $k\in\{2,3,4\}$, and $\alpha(k,\mathcal{B}_{3})=Z$ for all $k\in\{4,5,6\}$.
+Note that node $4$ has separate weights for the two bags it is part of.
+Figure 4: Bag generation and scoring of relation $c$.
+For solving the new classification task we score the candidate relations $c$ and $d$.
+The right part of Figure 4 illustrates the scoring of $c$.
+Due to the indistinguishable structure of the $c$-neighborhood for bags $\mathcal{B}_{2}$ and $\mathcal{B}_{3}$ the loss function 4 now reduces again to the constant $1/2$, as in the scoring of relation $a$.
+H Refer to caption
+Figure 5: Scoring relation $d$.
+For scoring the relation $d$ we obtain as the loss function 4
+$\sigma\left(Z\Theta^{T}\begin{bmatrix}1\\ 0\end{bmatrix}(w_{12}-w_{10}-w_{11})+Z\Theta^{T}\begin{bmatrix}0\\ 1\end{bmatrix}(w_{10}+w_{11}-1)\right)$
 […]
-Figure 2: Top: First iteration scoring function.
-For simplicity, we assume the features of the patient nodes to be uninformative, represented by a constant ’$T$’(rue) vector.
-Top Left: By simultaneous optimization of $w_{2},\ldots,w_{6}$ and $\Theta$ parameters, the loss for relation ’a’ can be brought arbitrarily close to 0 by choosing $\Theta$ so that $Z:=\Theta\cdot T\gg 0$.
-Top Right:The loss for relation "b" is constant $1/2$ for any choice of the parameters $w_{7},\Theta$.
-Bottom: Second iteration of the scoring function.
-With $Z$ as in first iteration we have $\alpha(2,\mathcal{B}_{2})=\alpha(3,\mathcal{B}_{2})=\alpha(5,\mathcal{B}_{3})=\alpha(6,\mathcal{B}_{3})=Z$ and $\alpha(4,\mathcal{B}_{2})=\alpha(4,\mathcal{B}_{3})=2Z$ (details in App.
-A.2).
-Bottom Left: With relation "c" optimal parameters give 0 loss.
-Bottom Right: Relation "d" obtains non-optimal loss (1/2).
+4.2.1 Time complexity analysis
+The purpose of the scoring function is to iteratively construct the meta-path that best classifies the target nodes, avoiding the need to explore all possible paths.
+A naive search algorithm that simply tests all possible paths would have a polynomial complexity of $O(|\mathcal{R}|^{L})$, where $|\mathcal{R}|$ is the number of relations in the graph, and $L$ is the maximum length of the meta-path we aim to find.
+By leveraging the scoring function, the complexity is reduced to linear $O(|\mathcal{R}|\cdot L)$.
+This improvement is achieved because, at each step, one relationship is added to the meta-path under construction, and the subsequent search builds upon it without reconsidering the other relationships scored in the same iteration.
 4.3 MPS-GNN
@@
 $h^{(l+1)}_{v}=\sigma\Big{(}W_{0}^{(l)}h^{(l)}_{v}+W_{neigh}^{(l)}\sum_{u\in\mathcal{N}_{v}^{r_{L-l}}}h^{(l)}_{u}+W_{1}^{(l)}h^{(0)}_{v}\Big{)}$ (7)
-where $\mathcal{N}_{v}^{r_{L-l}}$ are neighbours of node $v$ under relation $r_{L-l}$, $h^{(l)}_{v}$ is the embedding of node $v$ in layer $l$, $h^{(0)}_{v}=x_{v}$ is the feature vector of node $v$, while $W_{0}^{(l)}$, $W_{neigh}^{(l)}$ and $W_{1}^{(l)}$ are learnable parameter vectors.
-Note that the latter term $W_{1}^{(l)}h^{(0)}_{v}$, which is missing the original MP-GNN [29], represents a skip connection between the input and the $l+1$ layer.
-This allows the network to access the node attributes at each layer, which is essential for enabling the MPS-GNN to capture node features corresponding to the $\Theta\cdot x_{v}$ terms in the meta-path construction as shown in the ablation study in A.3).
-As in [29], the definition can be generalized to multiple meta-paths by concatenating the embeddings obtained from each of them.
+[…]
 4.4 The overall algorithm
-Algorithm 1 outlines the whole MPS-GNN procedure for the single meta-path case (in practice, a $K$ beam size is used and multiple meta-paths are learned).
-The algorithm takes as input a graph ${\cal G}$, the set of available relations ${\cal R}$, an initial set of binary node labels ${\cal Y}$ and a maximal meta-path length $L_{MAX}$.
+Algorithm 1 outlines the whole MPS-GNN procedure for the single meta-path case (in practice, a $K$ beam search is used and multiple meta-paths are learned).
+The algorithm takes as input a graph ${\cal G}$, the set of available relations ${\cal R}$, an initial set of binary node labels ${\cal Y}$ a maximal meta-path length $L_{MAX}$ and a stopping criteria value $\eta$.
 It initializes the targets ${\cal S}$ with a set of singletons (one per labelled node) and their alpha values (collectively indicated by $\mathcal{A}$) to 1.
 At each iteration, the scoring function identifies the relation minimizing Eq.
-3.
-The algorithm then constructs an MPS-GNN with the current (partial) meta-path, trains it on node labels, and evaluates it using the $F_{1}$ score (computed on a validation set, omitted for brevity), which reflects the meta-path’s performance when embedded in an MPS-GNN, as opposed to its potential informativeness measured by the scoring function.
+4 and appends it to the meta-path $mp$.
+If no relation improves the loss by a factor $\eta$, the algorithm stops and returns the meta-path $mp^{*}$ constructed so far.
+The algorithm then evaluates $mp$ by training MPS-GNN on node labels, and tests it using the $F_{1}$ score (computed on a validation set, omitted for brevity), which reflects the meta-path’s performance when embedded in an MPS-GNN, as opposed to its potential informativeness measured by the scoring function.
+For training MPS-GNN, node embeddings are updated using 7.
 Algorithm 1 MPS-GNN Learning
-procedure MPS-GNN(${\cal G},{\cal R},{\cal Y},L_{MAX}$)
+procedure learnMPS-GNN(${\cal G},{\cal R},{\cal Y},L_{MAX},{\eta}$)
      $mp^{*}\leftarrow[\,]$, $F_{1}^{*}\leftarrow 0$, ${\cal S}\leftarrow{\cal Y}$, $\mathcal{A}\leftarrow 1$
      while $|mp|<L_{MAX}$ do
-         $r^{*}\leftarrow\textsc{score}({\cal G},{\cal S},\mathcal{R},\mathcal{A})$
+         $r^{*}\leftarrow argmin_{r}L(r)$triangleright Eq. 4
+         if $\min_{r\in\mathcal{R}}L(r)\geq\eta L_{\text{init}}(r)$ then
+              return $mp^{*}$
+         end if
          $mp\leftarrow mp,r^{*}$
-[…]
+         $gnn\leftarrow\textsc{train}(\textsc{MPS-GNN}(mp),{\cal G},{\cal Y})$
+         $F_{1}\leftarrow\textsc{test}(gnn)$
          if $F_{1}>F_{1}^{*}$ then
@@
          end if
-         $\mathcal{A},{\cal S}\leftarrow\textsc{new-targets}({\cal S},r^{*})$
+         $\mathcal{A},{\cal S}\leftarrow\textsc{new-targets}({\cal S},r^{*})$triangleright Eq. 5, 6
      end while
@@
 end procedure
-The algorithm keeps track of the best meta-path prefix together to its $F_{1}$ score, and updates the bags and $\alpha$ values as specified in Equations 5 and 6 for the next relation scoring round.
-The algorithm ends when the maximum meta-path length $L_{MAX}$ is reached.
+The algorithm keeps track of the best meta-path prefix together to its $F_{1}$ score, and creates new target bags and $\alpha$ values as specified in Equations 5 and 6 for the next relation scoring round.
+The algorithm ends when the maximum meta-path length $L_{MAX}$ is reached The algorithm is described for simplicity in the context of a single meta-path and without the additional stopping criteria described in section 4.2.
+In practice, however, the implementation employs a beam search over the meta-path space, selecting the top $K$ best-scoring meta-paths concatenating their embedding for the final node representation as detailed at the end of section 4.3.
 MPS-GNN scales linearly in the number of relations and nodes, thanks to its incremental construction of meta-paths.
+The value chosen for $L_{\text{MAX}}$ in the experiment was 4, while $\eta$ was set to 0.7.
+The method is however quite robust to variations in $\eta$, as the identified meta-paths remained unchanged for a wide range of values in our experiments.
 4.5 MPS-GNN is a self-explainable model
-By relying on meta-paths for its predictions, MPS-GNN is a self-explainable model.
+Self-explainable GNNs (Kakkad et al., 2023; Christiansen et al., 2023) are a class of GNN models that aims to achieve explainability by-design.
+At a high level, these models can be seen as composed of two modules: a detector that extracts a class-discriminative subgraph, and a classifier that outputs a prediction based on the extracted subgraph.
+By relying on meta-paths for its predictions, MPS-GNN is a self-explainable GNN model.
 The scoring function serves as the detector, identifying relevant meta-paths, while the network built using them acts as the classifier.
@@
 An analysis of the faithfulness of MPS-GNN’s explanations is provided in Section 5.3.
+While most approaches for GNN explainability focus on the topological aspect, by identifying (hard or soft) subgraphs as explanations, determining which node features contribute to the prediction is also relevant from an interpretability perspective.
+While interpretability at the node feature level could be encouraged by introducing sparsifying norms for the learnable parameter vectors in Eq.
+7, guaranteeing a fully transparent processing of node features in the layerwise node embedding computation is beyond the scope of current GNN-based architectures.
+4.6 Comparison with MP-GNN
+The novelty of our approach compared to MP-GNN lies in our model’s ability to learn meta-paths that are relevant to the target node class, not merely based on their existence but on statistical measures related to their occurrences.
+In practice the key difference between MPS-GNN and MP-GNN lies in the way in which relations are scored.
+Specifically MP-GNN predicts the class label $y$ of a node $v$, in the first iteration, with $\tilde{y}_{v}^{r}=\Theta^{T}x_{v}\max_{u\in{\cal N}_{v}^{r}}w_{u}$ where the $max$ aggregation of the neighbours is used, indicating that a candidate relation $r$ is informative for the label of a node $v$ if at least one of the neighbors ${\cal N}_{v}^{r}$ of $v$ according to $r$ belongs to the ground-truth meta-path, and $v$ has the right features.
+To leverage meta-path occurrences, MPS-GNN employs a $sum$ aggregation strategy, as detailed in Eq.
+2, enabling the counting of their occurrences.
+In subsequent iterations, MP-GNN also employs a bag creation process, where positive bags are formed by including neighbors of positive nodes that are predicted to be positive at least once across multiple prediction procedures.
+This formulation ensures that the node responsible for the positive label, and thus aligned with the correct meta-path, remains in a positive bag.
+Unlike this approach, however, we must ensure that all neighbor nodes involved in multiple occurrences of the $correct$ meta-path are included in a positive bag.
+As detailed in eq.
+5, the neighbors of positive nodes are therefore placed in a positive bag without the need of any additional prediction step.
+The enhancements introduced in MPS-GNN, compared to MP-GNN, enable it to handle predictions over relational databases where the class label may depend on statistical measures derived from meta-path occurrences.
 5 Experiments
@@
 We compared MPS-GNN with approaches that don’t require predefined meta-paths, handle numerous relations, and incorporate node features in learning.
-The identified competitors include: MLP, to test the sufficiency of target node features alone; GCN [43], a baseline non-relational model; RGCN [44], extending GCN for multi-relational graphs, with distinct parameters for each edge type; HGN [45], a heterogeneous GNN model extending GAT for multiple relations; GTN [46], which transforms input graphs into different meta-path graphs where node representations are learned; Fast-GTN [47], an optimized GTN variant; R-HGNN [48], a relation-aware GNN using cross-relation message passing; and MP-GNN [29], the original meta-path GNN supporting only existentially quantified meta-paths.
+The identified competitors include: MLP, to test the sufficiency of target node features alone; GCN (Kipf & Welling, 2016), a baseline non-relational model; RGCN (Schlichtkrull et al., 2017), extending GCN for multi-relational graphs, with distinct parameters for each edge type; HGN (Lv et al., 2021a), a heterogeneous GNN model extending GAT for multiple relations; GTN (Yun et al., 2019a), which transforms input graphs into different meta-path graphs where node representations are learned; Fast-GTN (Yun et al., 2022b), an optimized GTN variant; R-HGNN (Yu et al., 2021), a relation-aware GNN using cross-relation message passing; and MP-GNN (Ferrini et al., 2024), the original meta-path GNN supporting only existentially quantified meta-paths.
 We implemented our model using PyTorch Geometric, and used the competitors’ code from their respective papers for comparison.
 For training MPS-GNN, we used a 70/20/10 split for training, validation, and testing, respectively, and reported the test results for the model selected based on its validation performance.
-For the sake of comparison with [29], we set the maximum meta-path length to 4 and the beam size to 3.
+For the sake of comparison with Ferrini et al.
+(2024), we set the maximum meta-path length to 4 and the beam size to 3.
 We employed $F_{1}$ as evaluation metric to account for the unbalancing in many of the datasets.
-The code is freely available at  [^1] https://github.com/francescoferrini/MPS-GNN
-5.1 Q1: MPS-GNN consistently identifies the correct meta-path in synthetic scenarios
+The code is freely available at  [^2] https://github.com/francescoferrini/MPS-GNN.
+Hyperparameters of competitors and MPS-GNN can be found in Table 8 in the Appendix.
+5.1 Q1: MPS-GNN consistently identifies the correct meta-path in count based synthetic scenarios
 In order to address the first research question, we designed a sequence of synthetic node classification scenarios where the correct structure to be learnt is known.
 In each scenario, a node is labelled as positive if it is the starting point of at least $c$ occurrences of a given meta-path of length $l$, and negative otherwise.
-Crucially, existential quantification of meta-paths (as modelled by MP-GNN [29]) is insufficient here, as nodes which are starting points of less than $c$ meta-path occurrences are labelled as negatives.
+Crucially, existential quantification of meta-paths (as modelled by MP-GNN (Ferrini et al., 2024)) is insufficient here, as nodes which are starting points of less than $c$ meta-path occurrences are labelled as negatives.
 We designed scenarios of increasing complexity by changing the length of the ground-truth meta-path $l$, the number of occurrences $c$, and the overall number of relations $r$ in the dataset.
 See Figure 6 for the statistics of the different scenarios (left), and for a prototypical example for $l=2$ and $c=3$ (right).
+Correct Pattern
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+S1 S2 S3 S4 S5 S6 S7 S8
+$|R|$ 5 5 5 5 10 10 10 10
+$c$ 2 2 3 4 2 2 3 4
+$l$ 2 3 3 2 2 3 3 2
 […]
-$|\mathcal{R}|$ $c$ $l$
-S1 5 2 2
-S2 5 2 3
-S3 5 3 3
-S4 5 4 2
-S5 10 2 2
-S6 10 2 3
-S7 10 3 3
-S8 10 4 2
 Table 1 shows the $F_{1}$ score of each model for an increasing complexity of the classification scenario.
@@
 Conversely, MPS-GNN manages to achieve nearly-optimal performance in all scenarios, substantially outperforming all existing strategies [^3] The residual error for MPS-GNN is due to the fact that despite relying on the correct meta-path, it occasionally leverages spurious instances where the relation sequence is correct but (some of) the node features are not..
+Note that the lookahead capabilities of the scoring function are crucial to the effectiveness of MPS-GNN.
+Appendix A.7 shows how replacing the scoring function with a simple greedy approach leads to failure in learning the correct meta-path.
 These results allow us to answer the first research question in the affirmative.
-Table 1: $F_{1}$ scores on the synthetic datasets. Best results are boldfaced. Results are averaged over 5 runs. The full table including standard deviations can be found in Appendix A.4
+Table 1: $F_{1}$ metric with standard deviations for synthetic datasets
 S1 S2 S3 S4 S5 S6 S7 S8
-MLP 0.46 0.44 0.48 0.47 0.44 0.51 0.45 0.47
-GCN 0.46 0.46 0.48 0.52 0.44 0.48 0.46 0.48
-RGCN 0.78 0.87 0.86 0.81 0.86 0.77 0.91 0.79
-R-HGNN 0.50 0.44 0.47 0.47 0.53 0.48 0.46 0.48
-HGN 0.45 0.46 0.50 0.46 0.46 0.48 0.45 0.40
-GTN 0.46 0.52 0.49 0.48 0.44 0.47 0.49 0.47
-Fast-GTN 0.46 0.48 0.51 0.49 0.44 0.47 0.53 0.47
-MP-GNN 0.84 0.82 0.85 0.95 0.89 0.79 0.84 0.71
-MPS-GNN 0.98 0.98 0.99 0.98 0.99 0.93 0.94 0.95
+MLP 0.46(±${\scriptstyle(\pm 0.00)}$ 0.44(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$ 0.47(±${\scriptstyle(\pm 0.00)}$ 0.44(±${\scriptstyle(\pm 0.00)}$ 0.51(±${\scriptstyle(\pm 0.00)}$ 0.45(±${\scriptstyle(\pm 0.00)}$ 0.47(±${\scriptstyle(\pm 0.00)}$
+GCN 0.46(±${\scriptstyle(\pm 0.00)}$ 0.46(±${\scriptstyle(\pm 0.02)}$ 0.48(±${\scriptstyle(\pm 0.03)}$ 0.52(±${\scriptstyle(\pm 0.05)}$ 0.44(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$ 0.46(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$
+RGCN 0.78(±${\scriptstyle(\pm 0.02)}$ 0.87(±${\scriptstyle(\pm 0.03)}$ 0.86(±${\scriptstyle(\pm 0.03)}$ 0.81(±${\scriptstyle(\pm 0.02)}$ 0.86(±${\scriptstyle(\pm 0.03)}$ 0.77(±${\scriptstyle(\pm 0.01)}$ 0.91(±${\scriptstyle(\pm 0.00)}$ 0.79(±${\scriptstyle(\pm 0.01)}$
+R-HGNN 0.50(±${\scriptstyle(\pm 0.00)}$ 0.44(±${\scriptstyle(\pm 0.03)}$ 0.47(±${\scriptstyle(\pm 0.01)}$ 0.47(±${\scriptstyle(\pm 0.04)}$ 0.53(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.01)}$ 0.46(±${\scriptstyle(\pm 0.02)}$ 0.48(±${\scriptstyle(\pm 0.02)}$
+HGN 0.45(±${\scriptstyle(\pm 0.00)}$ 0.46(±${\scriptstyle(\pm 0.00)}$ 0.50(±${\scriptstyle(\pm 0.03)}$ 0.46(±${\scriptstyle(\pm 0.00)}$ 0.46(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$ 0.45(±${\scriptstyle(\pm 0.03)}$ 0.40(±${\scriptstyle(\pm 0.12)}$
+GTN 0.46(±${\scriptstyle(\pm 0.00)}$ 0.52(±${\scriptstyle(\pm 0.00)}$ 0.49(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$ 0.44(±${\scriptstyle(\pm 0.00)}$ 0.47(±${\scriptstyle(\pm 0.00)}$ 0.49(±${\scriptstyle(\pm 0.00)}$ 0.47(±${\scriptstyle(\pm 0.00)}$
+Fast-GTN 0.46(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$ 0.51(±${\scriptstyle(\pm 0.00)}$ 0.49(±${\scriptstyle(\pm 0.00)}$ 0.44(±${\scriptstyle(\pm 0.00)}$ 0.46(±${\scriptstyle(\pm 0.00)}$ 0.53(±${\scriptstyle(\pm 0.00)}$ 0.47(±${\scriptstyle(\pm 0.00)}$
+MPGNN 0.84(±${\scriptstyle(\pm 0.09)}$ 0.82(±${\scriptstyle(\pm 0.13)}$ 0.85(±${\scriptstyle(\pm 0.10)}$ 0.95(±${\scriptstyle(\pm 0.02)}$ 0.89(±${\scriptstyle(\pm 0.06)}$ 0.79(±${\scriptstyle(\pm 0.03)}$ 0.84(±${\scriptstyle(\pm 0.06)}$ 0.71(±${\scriptstyle(\pm 0.21)}$
+MPS-GNN 0.98(±${\scriptstyle(\pm 0.00)}$ 0.98(±${\scriptstyle(\pm 0.01)}$ 0.99(±${\scriptstyle(\pm 0.10)}$ 0.98(±${\scriptstyle(\pm 0.00)}$ 0.99(±${\scriptstyle(\pm 0.00)}$ 0.93(±${\scriptstyle(\pm 0.10)}$ 0.94(±${\scriptstyle(\pm 0.00)}$ 0.95(±${\scriptstyle(\pm 0.00)}$
 5.2 Q2: MPS-GNN surpasses competitors in real world databases, learning relevant meta-paths
 Our approach is particularly useful for predictive tasks in relational databases with multiple tables, where features for a target entity may involve statistics from related tables.
-To address the second research question, we thus focused on three relational databases with many tables: EICU, a medical database with 31 tables, where we predict patient stay duration in the eICU; MONDIAL, a geographic database where the task is predicting whether a country’s religion is Christian; and ErgastF1, containing Formula 1 data, where the task is predicting the winner of a race.
+To address the second research question, we thus focused on three relational databases with many tables: EICU, a medical database with 31 tables, where we predict patient stay duration in the eICU, modeled as binary node classification by thresholding duration at 20 hours to achieve two balanced classes.; MONDIAL, a geographic database where the task is predicting whether a country’s religion is Christian; and ErgastF1, containing Formula 1 data, where the task is predicting the winner of a race in a binary classification task where target nodes are represented by a combination of race and pilot.
 The databases were transformed into graphs as explained in Section 3; for disconnected components, we enhanced connectivity by clustering rows of auxiliary tables.
-Additional details for the dataset and the procedure are in the Appendix.
+Additional details for the datasets and the procedure are in the Appendix A.3.
 Results
+Table 2: $F_{1}$-score with standard deviations of our method and competitors on real-world datasets.
+EICU MONDIAL ErgastF1
+MLP 0.53(±${\scriptstyle(\pm 0.02)}$ 0.52(±${\scriptstyle(\pm 0.00)}$ 0.50(±${\scriptstyle(\pm 0.00)}$
+GCN 0.89(±${\scriptstyle(\pm 0.00)}$ 0.60(±${\scriptstyle(\pm 0.02)}$ 0.50(±${\scriptstyle(\pm 0.01)}$
+RGCN 0.70(±${\scriptstyle(\pm 0.00)}$ 0.53(±${\scriptstyle(\pm 0.08)}$ 0.57(±${\scriptstyle(\pm 0.01)}$
+R-HGNN 0.61(±${\scriptstyle(\pm 0.00)}$ 0.61(±${\scriptstyle(\pm 0.01)}$ 0.72(±${\scriptstyle(\pm 0.02)}$
+HGN 0.75(±${\scriptstyle(\pm 0.00)}$ 0.72(±${\scriptstyle(\pm 0.01)}$ 0.70(±${\scriptstyle(\pm 0.04)}$
+GTN 0.56(±${\scriptstyle(\pm 0.02)}$ 0.38(±${\scriptstyle(\pm 0.01)}$ 0.60(±${\scriptstyle(\pm 0.01)}$
+Fast-GTN 0.46(±${\scriptstyle(\pm 0.03)}$ 0.39(±${\scriptstyle(\pm 0.04)}$ 0.60(±${\scriptstyle(\pm 0.03)}$
+MP-GNN 0.87(±${\scriptstyle(\pm 0.02)}$ 0.36(±${\scriptstyle(\pm 0.06)}$ 0.71(±${\scriptstyle(\pm 0.01)}$
+MPS-GNN 0.92(±${\scriptstyle(\pm 0.01)}$ 0.74(±${\scriptstyle(\pm 0.01)}$ 0.83(±${\scriptstyle(\pm 0.02)}$
 Table 2 presents the $F_{1}$ scores of MPS-GNN and its competitors across three real-world databases, averaged over 5 runs with different seeds.
@@
 Heterogeneous GNN methods also struggle with these datasets, especially MONDIAL, where most approaches fail to outperform a simple MLP, and only one (HGN) manages to substantially outperform the non-heterogeneous baseline (GCN).
-EICU MONDIAL ErgastF1
-MLP 0.53 0.52 0.50
-GCN 0.89 0.60 0.50
-RGCN 0.70 0.53 0.57
-R-HGNN 0.61 0.61 0.72
-HGN 0.75 0.72 0.70
-GTN 0.56 0.38 0.60
-Fast-GTN 0.46 0.39 0.60
-MP-GNN 0.87 0.36 0.71
-MPS-GNN 0.92 0.74 0.83
-Table 2: $F_{1}$-score of our method and competitors on the three real-world datasets. Full table with standard deviations can be found in Appendix A.6
-MP-GNN does not provide the performance boost that was observed when applied to knowledge graphs [29], confirming our intuition that existential quantification of meta-path is insufficient when dealing with relational databases.
+MP-GNN does not provide the performance boost that was observed when applied to knowledge graphs (Ferrini et al., 2024), confirming our intuition that existential quantification of meta-path is insufficient when dealing with relational databases.
 On the other hand, MPS-GNN manages to substantially outperform all competitors, thanks to its ability to identify meta-paths that are informative thanks to the statistics that can be computed over their realizations, as shown in the following.
@@
 Finally, in ErgastF1 the winner is predicted via meta-paths collecting information about the constructor (top) and driver (top) standings.
-Finally, in Appendix A.11 we present an experimental evaluation where MPS-GNN is adapted to deal with temporal databases and tasks [31], showing how it outperforms its competitors also in this context.
+Finally, in Appendix A.8 we present an experimental evaluation where MPS-GNN is adapted to deal with temporal databases and tasks (Robinson et al., 2024), showing how it outperforms its competitors also in this context.
 Summing up, these results enable us to confidently answer Q2 in the affirmative.
@@
 Indeed, the computational graph of MPS-GNN consists solely of the subgraph containing the occurrences of the identified meta-paths.
-Necessity, on the other hand, should be experimentally evaluated by altering the computational graph of MPS-GNN.
-Table 3: Changes to $F_{1}$ and necessity when removing 25%, 50%, and 75% of the learned meta-path occurrences for the real-world datasets.
-$\mathbf{F_{1}}$ Necessity
-Removed (%) 0 25 50 75 0 25 50 75
-EICU 0.92 0.82 0.67 0.56 0 0.16 0.18 0.24
-MONDIAL 0.74 0.61 0.48 0.43 0 0.13 0.29 0.32
-ErgastF1 0.814 0.65 0.61 0.58 0 0.08 0.16 0.26
-Table 3 illustrates the effects of removing 25%, 50%, and 75% of the meta-path occurrences in terms of changes in $F_{1}$ and necessity between original and modified graphs [49].
-For a visual representation of this table and for the definition of necessity, see Appendix A.7.
-In all datasets, there’s a noticeable decline in $F_{1}$ performance and a steep increase in probability difference, suggesting that the learned meta-paths are also necessary.
-These results clearly indicate the faithfulness of the explanations of our self-explainable method.
+Necessity, on the other hand, is calculated as a distance metric, measuring the difference in predicted probabilities between the original predictions and those obtained after masking parts of the explanation (i.e.
+deleting some instances of meta-paths).
+Defined as $\textsc{Nec}=\frac{1}{N}{\sum}_{v=1}^{N}(p_{v}(\mathcal{G})-p_{v}(\mathcal{G}^{\prime}))$ where $v$ is a target node, $\mathcal{G}$ is the original graph and $p_{v}$ denotes the probability associated with the predicted class.
+$\mathcal{G^{\prime}}$ is obtained by removing certain meta-path occurrences (i.e.
+randomly removing some branches of the identified meta-paths) and $p_{v}(\mathcal{G^{\prime}})$ is the probability associated at the class predicted with $p_{v}$.
+Since MPS-GNN utilizes just the explanation for making prediction one should expect that removing some part of the explanation, has as effect a decrease in prediction $F_{1}$.
+Figure 8: Changes to $F_{1}$ and posterior probability difference (necessity) when removing 25%, 50%, and 75% of the learned meta-path occurrences for the real-world datasets with MPS-GNN, dashed line, and MP-GNN, solid line.
+Figure 8 illustrates the effects of removing 25%, 50%, and 75% of the meta-path occurrences in terms of changes in $F_{1}$ and necessity between original and modified graphs (Azzolin et al., 2025).
+In addition to the results for MPS-GNN (dashed line), the figure includes the results for MP-GNN (solid line), which is also a self-explainable GNN model according to the reasoning in Section 4.5.
+In all datasets, there’s a noticeable decline in $F_{1}$ performance for MPS-GNN and a steep increase in probability difference, suggesting that the learned meta-paths are also necessary.
+These results clearly indicate the faithfulness of the explanations of MPS-GNN.
+MP-GNN has a similar behaviour, albeit with lower $F_{1}$ with respect to MPS-GNN because of its lower expressivity.
+The only exception is the MONDIAL dataset, where MP-GNN fails to learn any relevant pattern, resulting in a very low $F_{1}$ score that remains approximately constant when removing meta-path occurrences.
 6 Conclusion
 We introduced a novel approach to identify relevant meta-paths of relations for node classification tasks in heterogeneous graphs with a potentially large number of different relations, notably graphs derived from relational databases.
-Compared to earlier work, our approach does not require any user supervision, and it can learn meta-paths for predictive features defined by aggregate statistics over meta-path occurrences.
-Experiments clearly demonstrate advantages over alternative approaches in terms of accuracy, sparsity, and explainability.
-The approach was presented for binary classification, but it can be straightforwardly extended to multiclass classification using standard one-vs-all strategies.
-Extending it to regression tasks is an interesting direction for further research.
+Compared to earlier work, our approach does not require user supervision and learns meta-paths for predictive features defined by aggregate statistics over meta-path occurrences.
+The explainability of our method is particularly beneficial for sensitive domains like medical or financial data, where it helps address fairness concerns by providing insights into predictions.
+Experiments demonstrate advantages in accuracy and explainability.
+Limitations and future works
+At this stage, our method is tailored for binary node classification but can be extended to multiclass classification using standard one-vs-all strategies.
+However, this approach is inefficient, as it requires repeating the entire process for each class.
+The underlying principles of our local greedy meta-path construction based on scoring potential informativeness of relations directly applies also in the multiclass case.
+What would be needed in a multiclass adaptation of the approach is to replace the scalar functions (1),(2), and scalar node weights $w_{u}$, with vector-valued versions, and to modify the loss function (3) accordingly.
+Similarly, node regression is currently not supported and would require a modification to the way relations are scored.
+Another limitation of the scoring function is its reliance on a well-connected graph structure; when target nodes have neighbors that are not connected to other target nodes, the scoring function requires a preprocessing step to create supernodes within the neighborhood, as implemented in certain real-world scenarios.
+An interesting direction for future work involves incorporating temporal information into the process of learning optimal meta-paths.
+By doing so, the model would be capable of accounting for the temporal dimension, enabling it to better capture time-dependent relationships and dynamics within the graph.
+7 Acknowledgments
+Funded by the European Union.
+Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA).
+Neither the European Union nor the granting authority can be held responsible for them.
+Grant Agreement no.
+101120763 - TANGO.
+AL acknowledges the support of the MUR PNRR project FAIR - Future AI Research (PE00000013) funded by the NextGenerationEU.
 Appendix A Appendix
-A.1 Comparison with MP-GNN
-The novelty of our approach compared to MP-GNN lies in our model’s ability to learn meta-paths that are relevant to the target node class, not merely based on their existence but on statistical measures related to their occurrences.
-As illustrated in the example in Figure 1, MP-GNN would not be able to distinguish between the classes of the two patient nodes, as both have a single occurrence of the correct meta-path.
-This is due to how the function that iteratively constructs meta-paths in MP-GNN works.
-Specifically, under the existential quantification assumption, a candidate relation $r$ is informative for the label of a node $v$ f at least one of the neighbors ${\cal N}_{v}^{r}$ of $v$ according to $r$ belongs to the ground-truth meta-path, and $v$ has the right features.
-To achieve this, MP-GNN selects the optimal relation by identifying the connected node that maximizes information for class label prediction.
-This represents a significant limitation for MP-GNN, as it cannot differentiate between node classes that are determined by aggregate statistics (e.g.
-counts) of multiple occurrences of meta-paths.
-A.2 Second iteration loss computation
-In this Section we show the loss computation and minimization for relation "c" of the example in Figure 2.
-Note that $Z$ refers to the learned parameters related to the features of the target nodes in the previous iteration as shown in Figure 2 and $\binom{1}{0}$ and $\binom{0}{1}$ represent the features of $prescription$ nodes.
+A.1 Second iteration loss computation
+In this Section we show the loss computation and minimization for relation "d" of the example in Figure 5.
+Note that $Z$ refers to the learned parameters related to the features of the target nodes in the previous iteration with relation $b$ (Figure 3(b)) and $\begin{bmatrix}1\\ 0\end{bmatrix}$ and $\begin{bmatrix}0\\ 1\end{bmatrix}$ represent the features of $prescription$ nodes.
 $Z$ $\gg 0$
-$F(\mathcal{B}_{2},c)$ $=Z\Theta\binom{0}{1}+2Z\Theta\binom{1}{0}(w_{8}+w_{9})+Z\Theta\binom{1}{0}(w_{10}+w_{11})$
-$F(\mathcal{B}_{3},c)$ $=2Z\Theta\binom{1}{0}(w_{8}+w_{9})+Z\Theta\binom{0}{1}(w_{10}+w_{11})+Z\Theta\binom{1}{0}(w_{12})$
-$L(c)$ $=\min_{\Theta,w}\sigma\left(F(\mathcal{B}_{3},c)-F(\mathcal{B}_{2},c)\right)$
-$=\min_{\Theta,w}\sigma\left(Z\Theta\binom{1}{0}(2w_{8}+2w_{9}+w_{12}-2w_{8}-2w_{9}-w_{10}-w_{11})\right.$
-$\quad\left.+Z\Theta\binom{0}{1}(w_{10}+w_{11}-1)\right)$
-$=\min_{\Theta,w}\sigma\left(Z\Theta\binom{1}{0}(w_{12}-w_{10}-w_{11})+Z\Theta\binom{0}{1}(w_{10}+w_{11}-1)\right)$
-$\Theta\binom{1}{0}$ $\gg 0;\quad\Theta\binom{0}{1}=0;\quad w_{12}=0;\quad w_{10},w_{11}=1$
-A.3 Ablation study
+$F(\mathcal{B}_{2},d)$ $=Z\Theta^{T}\begin{bmatrix}0\\ 1\end{bmatrix}+2Z\Theta^{T}\begin{bmatrix}1\\ 0\end{bmatrix}(w_{8}+w_{9})+Z\Theta^{T}\begin{bmatrix}1\\ 0\end{bmatrix}(w_{10}+w_{11})$
+$F(\mathcal{B}_{3},d)$ $=2Z\Theta^{T}\begin{bmatrix}1\\ 0\end{bmatrix}(w_{8}+w_{9})+Z\Theta^{T}\begin{bmatrix}0\\ 1\end{bmatrix}(w_{10}+w_{11})+Z\Theta^{T}\begin{bmatrix}1\\ 0\end{bmatrix}(w_{12})$
+$L(d)$ $=\min_{\Theta,w}\sigma\left(F(\mathcal{B}_{3},d)-F(\mathcal{B}_{2},d)\right)$
+$=\min_{\Theta,w}\sigma\left(Z\Theta^{T}\begin{bmatrix}1\\ 0\end{bmatrix}(2w_{8}+2w_{9}+w_{12}-2w_{8}-2w_{9}-w_{10}-w_{11})\right.$
+$\quad\left.+Z\Theta^{T}\begin{bmatrix}0\\ 1\end{bmatrix}(w_{10}+w_{11}-1)\right)$
+$=\min_{\Theta,w}\sigma\left(Z\Theta^{T}\begin{bmatrix}1\\ 0\end{bmatrix}(w_{12}-w_{10}-w_{11})+Z\Theta^{T}\begin{bmatrix}0\\ 1\end{bmatrix}(w_{10}+w_{11}-1)\right)$
+$\Theta^{T}\begin{bmatrix}1\\ 0\end{bmatrix}$ $\gg 0;\quad\Theta^{T}\begin{bmatrix}0\\ 1\end{bmatrix}=0;\quad w_{12}=0;\quad w_{10},w_{11}=1$
+A.2 Ablation study
 In this section, we highlight the importance of the skip connection $W_{1}^{(l)}h_{v}^{(0)}$ considered in equation 7.
-Table 4 and 5 demonstrate respectively the significance of skip connections in the synthetic and real-world settings.
+Table 3 and 4 demonstrate respectively the significance of skip connections in the synthetic and real-world settings.
 Without considering the initial target node features, the $F_{1}$ score drops significantly, underscoring the critical role these features play.
+Table 3: $F_{1}$ metric with standard deviations for synthetic datasets with MPS-GNN and MPS-GNN without using skip connection.
 S1 S2 S3 S4 S5 S6 S7 S8
@@
 […]
-Table 4: $F_{1}$ metric with standard deviations for synthetic datasets with MPS-GNN and MPS-GNN without using skip connection.
+Table 4: $F_{1}$ metric with standard deviations for real-world datasets with MPS-GNN and MPS-GNN without using skip connection.
 EICU MONDIAL ErgastF1
@@
 […]
-Table 5: $F_{1}$ metric with standard deviations for real-world datasets with MPS-GNN and MPS-GNN without using skip connection.
-A.4 Synthetic datasets
-In table 6 are presented $F_{1}$-score with standard deviations of MPS-GNN and competitors in the synthetic settings.
-S1 S2 S3 S4 S5 S6 S7 S8
-MLP 0.46(±${\scriptstyle(\pm 0.00)}$ 0.44(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$ 0.47(±${\scriptstyle(\pm 0.00)}$ 0.44(±${\scriptstyle(\pm 0.00)}$ 0.51(±${\scriptstyle(\pm 0.00)}$ 0.45(±${\scriptstyle(\pm 0.00)}$ 0.47(±${\scriptstyle(\pm 0.00)}$
-GCN 0.46(±${\scriptstyle(\pm 0.00)}$ 0.46(±${\scriptstyle(\pm 0.02)}$ 0.48(±${\scriptstyle(\pm 0.03)}$ 0.52(±${\scriptstyle(\pm 0.05)}$ 0.44(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$ 0.46(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$
-RGCN 0.78(±${\scriptstyle(\pm 0.02)}$ 0.87(±${\scriptstyle(\pm 0.03)}$ 0.86(±${\scriptstyle(\pm 0.03)}$ 0.81(±${\scriptstyle(\pm 0.02)}$ 0.86(±${\scriptstyle(\pm 0.03)}$ 0.77(±${\scriptstyle(\pm 0.01)}$ 0.91(±${\scriptstyle(\pm 0.00)}$ 0.79(±${\scriptstyle(\pm 0.01)}$
-R-HGNN 0.50(±${\scriptstyle(\pm 0.00)}$ 0.44(±${\scriptstyle(\pm 0.03)}$ 0.47(±${\scriptstyle(\pm 0.01)}$ 0.47(±${\scriptstyle(\pm 0.04)}$ 0.53(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.01)}$ 0.46(±${\scriptstyle(\pm 0.02)}$ 0.48(±${\scriptstyle(\pm 0.02)}$
-HGN 0.45(±${\scriptstyle(\pm 0.00)}$ 0.46(±${\scriptstyle(\pm 0.00)}$ 0.50(±${\scriptstyle(\pm 0.03)}$ 0.46(±${\scriptstyle(\pm 0.00)}$ 0.46(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$ 0.45(±${\scriptstyle(\pm 0.03)}$ 0.40(±${\scriptstyle(\pm 0.12)}$
-GTN 0.46(±${\scriptstyle(\pm 0.00)}$ 0.52(±${\scriptstyle(\pm 0.00)}$ 0.49(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$ 0.44(±${\scriptstyle(\pm 0.00)}$ 0.47(±${\scriptstyle(\pm 0.00)}$ 0.49(±${\scriptstyle(\pm 0.00)}$ 0.47(±${\scriptstyle(\pm 0.00)}$
-Fast-GTN 0.46(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$ 0.51(±${\scriptstyle(\pm 0.00)}$ 0.49(±${\scriptstyle(\pm 0.00)}$ 0.44(±${\scriptstyle(\pm 0.00)}$ 0.46(±${\scriptstyle(\pm 0.00)}$ 0.53(±${\scriptstyle(\pm 0.00)}$ 0.47(±${\scriptstyle(\pm 0.00)}$
-MPGNN 0.84(±${\scriptstyle(\pm 0.09)}$ 0.82(±${\scriptstyle(\pm 0.13)}$ 0.85(±${\scriptstyle(\pm 0.10)}$ 0.95(±${\scriptstyle(\pm 0.02)}$ 0.89(±${\scriptstyle(\pm 0.06)}$ 0.79(±${\scriptstyle(\pm 0.03)}$ 0.84(±${\scriptstyle(\pm 0.06)}$ 0.71(±${\scriptstyle(\pm 0.21)}$
-MPS-GNN 0.98(±${\scriptstyle(\pm 0.00)}$ 0.98(±${\scriptstyle(\pm 0.01)}$ 0.99(±${\scriptstyle(\pm 0.10)}$ 0.98(±${\scriptstyle(\pm 0.00)}$ 0.99(±${\scriptstyle(\pm 0.00)}$ 0.93(±${\scriptstyle(\pm 0.10)}$ 0.94(±${\scriptstyle(\pm 0.00)}$ 0.95(±${\scriptstyle(\pm 0.00)}$
-Table 6: $F_{1}$ metric with standard deviations for synthetic datasets
-A.5 Real world setting
+A.3 Real world setting
 In our real-world scenario, we utilized three relational databases, which are detailed below.
@@
 EICU
-Medical database with 31 tables (node types) [^3] https://eicu-crd.mit.edu.
+Medical database with 31 tables (node types) [^4] https://eicu-crd.mit.edu from Johnson et al.
+(2021).
 The task is predicting the duration of a patient’s stay in the eICU after admission, modeled as binary node classification by thresholding duration at 20 hours to achieve two balanced classes.
@@
 MONDIAL
-Database  [^4] https://relational-data.org/dataset/Mondial containing data from multiple geographical web data sources [50].
+Database  [^5] https://relational-data.org/dataset/Mondial containing data from multiple geographical web data sources (May, 1999).
 We predict the religion of a country as Christian (positive) with 114 instances vs.
 all other religions with 71 instances.
-In this dataset, clustering of tables is done using DBSCAN [51] clustering algorithm.
+In this dataset, clustering of tables is done using DBSCAN (Ester et al., 1996) clustering algorithm.
 Table 7 shows the resulting number of clusters for each table of the original database. Clustering is computed using DBSCAN algorithm.
@@
 The objective is to predict the winner of a race using the data available before the race starts, such as the list of participants and qualifying times, while the actual lap times during the race are not yet available.
+Table 5: Setting of real-world datasets.
+$|\mathcal{T}|$ and $|\mathcal{R}|$ refers respectively to the total number of tables in the original database and the number of relations in the graph used by the models.
+$Rows$ is the sum of all the rows of each specific database.
 Database $|\mathcal{T}|$ $|\mathcal{R}|$ $Rows$
@@
 ErgastF1 14 33 544056
-Table 7: Setting of real-world datasets.
-$|\mathcal{T}|$ and $|\mathcal{R}|$ refers respectively to the total number of tables in the original database and the number of relations in the graph used by the models.
-$Rows$ is the sum of all the rows of each specific database.
+Table 6: Tables from the EICU dataset.
+’Clustering Feature’ refers to the feature used to group the rows in each table.
+If not present, it means that the specific table does not have any feature for that purpose, so the DBSCAN algorithm is employed.
+$Clusters$ indicates the final number of nodes after the clustering step.
 Table name Clustering Feature $Rows$ $Clusters$
@@
 vitalperiodic - 1634960 200
-Table 8: Tables from the EICU dataset.
-’Clustering Feature’ refers to the feature used to group the rows in each table.
-If not present, it means that the specific table does not have any feature for that purpose, so the DBSCAN algorithm is employed.
-$Clusters$ indicates the final number of nodes after the clustering step.
+Table 7: Tables from the MONDIAL dataset. $Clusters$ indicates the final number of clusters after applying DBSCAN algorithm on the features of the specific table.
 Table name $Rows$ $Clusters$
@@
 mountain 241 40
-Table 9: Tables from the MONDIAL dataset. $Clusters$ indicates the final number of clusters after applying DBSCAN algorithm on the features of the specific table.
-A.6 Real world setting results
-EICU MONDIAL ErgastF1 rel-f1-dnf rel-f1-top3
-MLP 0.53(±${\scriptstyle(\pm 0.02)}$ 0.52(±${\scriptstyle(\pm 0.00)}$ 0.50(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$
-GCN 0.89(±${\scriptstyle(\pm 0.00)}$ 0.60(±${\scriptstyle(\pm 0.02)}$ 0.50(±${\scriptstyle(\pm 0.01)}$ 0.57(±${\scriptstyle(\pm 0.01)}$ 0.52(±${\scriptstyle(\pm 0.02)}$
-RGCN 0.70(±${\scriptstyle(\pm 0.00)}$ 0.53(±${\scriptstyle(\pm 0.08)}$ 0.57(±${\scriptstyle(\pm 0.01)}$ 0.44(±${\scriptstyle(\pm 0.02)}$ 0.54(±${\scriptstyle(\pm 0.02)}$
-R-HGNN 0.61(±${\scriptstyle(\pm 0.00)}$ 0.61(±${\scriptstyle(\pm 0.01)}$ 0.72(±${\scriptstyle(\pm 0.02)}$ 0.60(±${\scriptstyle(\pm 0.01)}$ 0.63(±${\scriptstyle(\pm 0.02)}$
-HGN 0.75(±${\scriptstyle(\pm 0.00)}$ 0.72(±${\scriptstyle(\pm 0.01)}$ 0.70(±${\scriptstyle(\pm 0.04)}$ 0.61(±${\scriptstyle(\pm 0.02)}$ 0.61(±${\scriptstyle(\pm 0.01)}$
-GTN 0.56(±${\scriptstyle(\pm 0.02)}$ 0.38(±${\scriptstyle(\pm 0.01)}$ 0.60(±${\scriptstyle(\pm 0.01)}$ 0.41(±${\scriptstyle(\pm 0.02)}$ 0.45(±${\scriptstyle(\pm 0.01)}$
-Fast-GTN 0.46(±${\scriptstyle(\pm 0.03)}$ 0.39(±${\scriptstyle(\pm 0.04)}$ 0.60(±${\scriptstyle(\pm 0.03)}$ 0.51(±${\scriptstyle(\pm 0.01)}$ 0.50(±${\scriptstyle(\pm 0.02)}$
-MP-GNN 0.87(±${\scriptstyle(\pm 0.02)}$ 0.36(±${\scriptstyle(\pm 0.06)}$ 0.71(±${\scriptstyle(\pm 0.01)}$ 0.54(±${\scriptstyle(\pm 0.02)}$ 0.52(±${\scriptstyle(\pm 0.02)}$
-MPS-GNN 0.92(±${\scriptstyle(\pm 0.01)}$ 0.74(±${\scriptstyle(\pm 0.01)}$ 0.83(±${\scriptstyle(\pm 0.02)}$ 0.62(±${\scriptstyle(\pm 0.02)}$ 0.65(±${\scriptstyle(\pm 0.01)}$
-Table 10: $F_{1}$-score with standard deviations of our method and competitors on real-world datasets.
-A.7 Necessity calculation
-To assess whether an explanation is faithful to the model, it is crucial to identify the elements that significantly influence the model’s performance.
-In other words, if any part of this explanation is removed, the prediction should deteriorate.
-This concept is known as Necessity, as it evaluates which components of the explanation are indispensable for accurate predictions.
-To quantify this, a distance metric is used, measuring the difference in predicted probabilities between the original predictions and those obtained after masking parts of the explanation (i.e.
-deleting some instances of meta-paths).
-This metric is calculated as:
-$\textsc{Nec}=\frac{1}{N}{\sum}_{v=1}^{N}(p_{v}(\mathcal{G})-p_{v}(\mathcal{G}^{\prime}))$ (8)
-where $v$ is a target node, $\mathcal{G}$ is the original graph and $p_{v}$ denotes the probability associated with the predicted class.
-$\mathcal{G^{\prime}}$ is obtained by removing certain meta-path occurrences and $p_{v}(\mathcal{G^{\prime}})$ is the probability associated at the class predicted with $p_{v}$.
-Since MPS-GNN utilizes just the explanation for making prediction one should expect that removing some part of the explanation, has as effect a decrease in prediction $F_{1}$.
-Figure 5 shows five plots, one for each dataset.
-As in Table 3 each graph illustrates the effects of removing 25%, 50%, and 75% of the meta-path occurrences in terms of changes in $F_{1}$ and necessity between original and modified graphs [49].
-Figure 5: Changes to $F_{1}$ and posterior probability difference (necessity) when removing 25%, 50%, and 75% of the learned meta-path occurrences for the real-world datasets.
-A.8 Number of parameters
+A.4 Details of competitors’ architectures
+In Table 8, we present the hyperparameters for all competitors and MPS-GNN on the real-world datasets.
+For the synthetic cases, the only difference lies in the # layers, which is adjusted to match the length of the correct meta-paths for the competitors.
+For the losses CE is cross-entropy and nll is negative log likelihood.
+Table 8: Hyperparameters of competitors and MPS-GNN for the real-world datasets.
+The optimizer is omitted from the table as it is Adam for all models.
+lr denotes the learning rate, wd represents the weight decay, and Patience indicates the early stopping patience (if applicable).
+Hyperparameters # layers Embedding dim. lr wd # epochs Patience Loss
+MLP 2 32 0.01 0.0005 500 50 nll
+GCN 2 32 0.01 0.0005 500 50 nll
+RGCN 2 32 0.01 0.0005 500 50 nll
+R-HGNN 2 64 0.001 0 200 50 CE
+HGN 2 64 0.0001 0.0005 300 30 nll
+GTN 2 64 0.01 0.001 200 - CE
+Fast-GTN 2 64 0.01 0.001 200 - CE
+MP-GNN 2 32 0.01 0.0005 500 50 nll
+MPS-GNN 2 32 0.01 0.0005 500 50 nll
+A.5 Number of parameters
 In Table 9, we present the total number of parameters required for evaluating the various models.
@@
 In contrast, HGN and R-HGNN exhibit an exponential increase in the number of parameters.
+Table 9: Number of parameters for each model across synthetic and real-world datasets.
 MLP GCN RGCN R-HGNN HGN GTN Fast-GTN MP-GNN MPS-GNN
@@
 ErgastF1 4356 198142 5422432 396543021 439021 2542354 11325242 23413 29538
-Table 11: Number of parameters for each model across synthetic and real-world datasets.
-A.9 Execution times
+A.6 Execution times
 In Table 10, we present the training times for each model across the individual datasets.
@@
 Overall, our approach is consistently neither the quickest nor the slowest, yet it reliably achieves the highest average F1 score across all settings.
+Table 10: Training times, in seconds, for each model across synthetic and real-world datasets.
 MLP GCN RGCN R-HGNN HGN GTN Fast-GTN MP-GNN MPS-GNN
@@
 ErgastF1 543 954 1491 2456 2245 3015 2945 2280 2421
-Table 12: Training times, in seconds, for each model across synthetic and real-world datasets.
-A.10 Scoring Function Lookahead Illustration
+A.7 Scoring Function Lookahead Illustration
 We report an additional experiment demonstrating the lookahead capabilities implemented in the scoring function of our method, and its advantage over a simplistic greedy approach.
@@
 Comparing with the $F_{1}$ score of the ground truth metapath we find that, indeed, starting the metapath with $r_{2}$ is a suboptimal choice, and that our scoring function correctly identifies the most informative relation to start the metapath with, even though this informativeness only is materialized after extension of the metapath with $r_{2}$.
+Table 11: Comparison: our metapath construction vs. simple greedy alternative.
 Iteration 1 $r_{2}$- Extensions Ground truth
@@
 $F_{1}$ 0.79 0.82 0.69 0.83 0.82 0.85 0.99
-Table 13: Comparison: our metapath construction vs. simple greedy alternative.
-A.11 Temporal experiment
-Recently, a novel benchmark, rel-bench [31], has been introduced.
+A.8 Temporal experiment
+Recently, a novel benchmark, rel-bench Robinson et al.
+(2024), has been introduced.
 This benchmark consists of multiple relational databases represented as temporal graphs.
@@
 In Section A.8.1, we provide the necessity calculations for these datasets.
+Table 12: Setting of temporal real-world dataset.
 Database $|\mathcal{T}|$ $|\mathcal{R}|$ $Rows$
 rel-f1 9 26 97606
-Table 14: Setting of temporal real-world dataset.
+Table 13: $F_{1}$-score with standard deviations of our method and competitors on two temporal datasets.
 rel-f1-dnf rel-f1-top3
@@
 MPS-GNN 0.62(±${\scriptstyle(\pm 0.02)}$ 0.65 (±${\scriptstyle(\pm 0.01)}$
-Table 15: $F_{1}$-score with standard deviations of our method and competitors on two temporal datasets.
-A.11.1 Necessity calculation in temporal tasks
+A.8.1 Necessity calculation in temporal tasks
 In this section, we present the necessity calculation for the temporal datasets.
-Specifically, as detailed in Appendix A.7, Table 16 and figure6 demonstrates that removing certain identified meta-paths results in decreased performance ($F_{1}$) and an increased necessity value.
+Specifically, as detailed in Section 5.3, Table 14 and figure9 demonstrates that removing certain identified meta-paths results in decreased performance ($F_{1}$) and an increased necessity value.
 This finding confirms that the learned meta-paths are essential for the prediction task in these datasets as well.
+Table 14: Changes to $F_{1}$ and necessity when removing 25%, 50%, and 75% of the learned meta-path occurrences for the real-world temporal tasks rel-f1-dnf and rel-f1-top3.
 $\mathbf{F_{1}}$ Necessity
@@
 rel-f1-top3 0.65 0.62 0.55 0.50 0 0.19 0.34 0.34
-Table 16: Changes to $F_{1}$ and necessity when removing 25%, 50%, and 75% of the learned meta-path occurrences for the real-world temporal tasks rel-f1-dnf and rel-f1-top3.
 Figure 9: Changes to $F_{1}$ and posterior probability difference (necessity) when removing 25%, 50%, and 75% of the learned meta-path occurrences for the real-world temporal tasks rel-f1-dnf and rel-f1-top3.
-A.12 Non-GNN models on real-world databases
+A.9 Non-GNN models on real-world databases
 For the MONDIAL and ErgastF1 databases, non-GNN methods have shown competitive performance in the past.
-For example, [52, 53] report $F_{1}$ scores of $0.78$ and $0.77$ on MONDIAL, $0.4$ and $0.3$ points higher than MPS-GNN.
+For example, Schulte et al.
+(2013); Bina et al.
+(2013) report $F_{1}$ scores of $0.78$ and $0.77$ on MONDIAL, $0.4$ and $0.3$ points higher than MPS-GNN.
 However, these results are achieved on a simplified version of the database with only 12 tables, requiring manual feature selection.
@@
 By comparison, MPS-GNN is designed to handle scenarios with a large number of relations effectively.
+A.10 Toy example with more complex features
+Figure 10: Toy example similar to Figure 1 with more complex features on prescription nodes.
+As depicted in the legend (right), the first two positions of the feature vector determine an exempt prescription ($[1,0]$) or a not exempt prescription ($[0,1]$).
+The last value determine the price of the prescription in dollar.
+The highlighted subgraph shows a prototypical counts-of-counts pattern characterising positive patients, namely having exempt prescriptions where the total cost is more then 100 dollars, each containing at least two medications
+In Figure 10, we present a scenario similar to the one depicted in Figure 1.
+The key difference lies in the prescription nodes, which now have more complex feature representations.
+Specifically, the last value of the feature vector corresponds to the price of each prescription.
+The first iteration of the scoring function follows the same process as in the previous toy example, as illustrated in Figures 3(a) and 3(b).
+The most notable aspect, however, is the second iteration, where the target nodes are prescription nodes, similar to Figure 5, where the model evaluates relation $d$ and  $c$.
+Here, we provide the computations involved in minimizing the loss.
+Notably, only relation $d$ allows the loss to be minimized to 0.
+Therefore, relation $d$ will be the one included in the identified meta-path
+Minimization relation $c$
+$Z$ $\gg 0$
+$F(\mathcal{B}_{2},c)$ $=2Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 60\end{bmatrix}+Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 70\end{bmatrix}(w_{13})+Z\Theta^{T}\begin{bmatrix}0\\ 1\\ 20\end{bmatrix}$
+$F(\mathcal{B}_{3},c)$ $=2Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 60\end{bmatrix}+Z\Theta^{T}\begin{bmatrix}0\\ 1\\ 20\end{bmatrix}+Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 70\end{bmatrix}(w_{13})$
+$L(c)$ $=\min_{\Theta,w}\sigma\left(F(\mathcal{B}_{3},d)-F(\mathcal{B}_{2},d)\right)$
+$=\min_{\Theta,w}\sigma\left(0\right)$
+$=\frac{1}{2}$
+Minimization relation $d$
+$Z$ $\gg 0$
+$F(\mathcal{B}_{2},d)$ $=2Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 60\end{bmatrix}(w_{8}+w_{9})+Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 70\end{bmatrix}(w_{10}+w_{11})+Z\Theta^{T}\begin{bmatrix}0\\ 1\\ 20\end{bmatrix}$
+$F(\mathcal{B}_{3},d)$ $=2Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 60\end{bmatrix}(w_{8}+w_{9})+Z\Theta^{T}\begin{bmatrix}0\\ 1\\ 20\end{bmatrix}(w_{10}+w_{11})+Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 70\end{bmatrix}(w_{12})$
+$L(d)$ $=\min_{\Theta,w}\sigma\left(F(\mathcal{B}_{3},d)-F(\mathcal{B}_{2},d)\right)$
+$=\min_{\Theta,w}\sigma\left(Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 70\end{bmatrix}(w_{12})+Z\Theta^{T}\begin{bmatrix}0\\ 1\\ 20\end{bmatrix}(w_{10}+w_{11})-Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 70\end{bmatrix}(w_{10}+w_{11})-Z\Theta^{T}\begin{bmatrix}0\\ 1\\ 20\end{bmatrix}\right)$
+$=\min_{\Theta,w}\sigma\left(Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 70\end{bmatrix}(w_{12})+Z\Theta^{T}\begin{bmatrix}0\\ 1\\ 20\end{bmatrix}(w_{10}+w_{11}-1)-Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 70\end{bmatrix}(w_{10}+w_{11})\right)$
+$\Theta^{T}\begin{bmatrix}0\\ 1\\ 20\end{bmatrix}$ $\gg 0;\quad w_{10},w_{11},w_{12}=0$